### v1

# PaperRegister: Boosting Flexible-grained Paper Search via Hierarchical Register Indexing

Abstract
Paper search is an important activity for researchers, typically involving using a query with description of a topic to find relevant papers. As research deepens, paper search requirements may become more flexible, sometimes involving specific details such as module configuration rather than being limited to coarse-grained topics. However, previous paper search systems are unable to meet these flexible-grained requirements, as these systems mainly collect paper abstracts to construct index of corpus, which lack detailed information to support retrieval by finer-grained queries. In this work, we propose PaperRegister, consisted of offline hierarchical indexing and online adaptive retrieval, transforming traditional abstract-based index into hierarchical index tree for paper search, thereby supporting queries at flexible granularity. Experiments on paper search tasks across a range of granularity demonstrate that PaperRegister achieves the state-of-the-art performance, and particularly excels in fine-grained scenarios, highlighting the good potential as an effective solution for flexible-grained paper search in real-world applications.
Introduction
Paper search is an important and almost everyday activity for researchers (Kuhlthau 1991; Ellis, Cox, and Hall 1993; Hemminger et al. 2007; Case and Given 2016). Typically, this process begins when the user submits a natural-language query describing a topic, then the retrieval system matches this query against the paper corpus and returns the subset of papers with the highest relevance (Wadden et al. 2020; Cohan et al. 2020; Ajith et al. 2024; He et al. 2025). As researchers delve more deeply into their work, the paper search requirements can become increasingly flexible. For example, the view of input query may refer to detailed module configuration or methodological operation, rather than being limited to the level of coarse-grained topic (Mysore et al. 2021; Kang et al. 2024; Wang et al. 2023; Zhang et al. 2025). Therefore, a paper search system supporting queries at flexible levels of granularity is of important value.
Unfortunately, existing paper search systems cannot effectively handle queries across flexible granularity, since they mainly use paper abstracts to construct index of corpus for retrieval (Zheng et al. 2020; Gao et al. 2023; Mackie, Chatterjee, and Dalton 2023; Lei et al. 2024). When a query’s view involves finer‐grained information that does not appear in abstract, they fails to retrieve relevant papers. As shown in Figure 1, the input query “Retrieve papers about jointly training encoder and generator via minimizing the negative marginal log-likelihood without document encoder” specifies a detailed training operation, expecting to find the relevant paper. However, the target cannot be successfully retrieved by traditional paper search systems via abstract-based index, because the abstract does not mention such detailed training operation. Therefore, building a powerful paper search system that supports queries at flexible granularity remains an important open challenge.
Figure 1: PaperRegister can support flexible-grained paper search through the hierarchical register, while traditional method fails due to abstract cannot contain required details.
Figure 2: The overview of PaperRegister, including hierarchical indexing and adaptive retrieval. In offline stage, PaperRegister constructs hierarchical index tree via fine-grained content extracting and bottom-up content aggregating based on hierarchical register schema. In online stage, PaperRegister first identify the views of input query and then conduct view-based matching.
To this end, we propose PaperRegister, which transforms traditional abstract-based index into hierarchical index tree for paper search, thereby supporting queries at flexible granularity. Specially, PaperRegister includes offline hierarchical indexing and online adaptive retrieval. In the offline stage, to construct a hierarchical index tree for corpus, PaperRegister uses a hierarchical register schema, consisted of information nodes at various granularity, in which each node path represents a kind of view. Based on this schema, PaperRegister first extracts fine‐grained contents for each paper, then aggregates contents bottom‐up, layer by layer, yielding hierarchical register of the paper. And then hierarchical index tree is constructed by merging the register of all papers. In the online stage, to adaptively retrieve via the hierarchical index tree, PaperRegister first employs a view recognizer to identify views of input query, which associates to suitable indexes from the hierarchical index tree, then matches query against these indexes to achieve precise paper search.
One key aspect of PaperRegister is the view recognizer, which must be both low latency and high accuracy. To this end, we construct the view recognizer by training a small-scale language model via hierarchical-reward reinforcement learning. Firstly, in order to alleviate the online latency, we adopt a language model with 0.6 billion parameter as base. Then, in order to equip the model with basic identifying capability, we perform supervised fine-tuning (SFT) on it, with query as input and the golden view as label. Furthermore, in order to further enhance accuracy of the view recognizer, considering the hierarchical dependency of views in the hierarchical register schema, we design a hierarchical reward as the reinforcement learning signal, which is calculated by closeness level between the predicted view and golden view in hierarchical register schema. And then we employ this reward on group relative policy optimization (GRPO)  (Shao et al. 2024) training to enhance the view recognizer.
In experiments, we compare PaperRegister with some common-used baselines on paper search tasks across various levels of granularity. The results show PaperRegister achieves state-of-the-art performance, with improvements being more pronounced as query granularity finer, confirming that PaperRegister is a robust solution for challenging flexible-grained paper search tasks. Additionally, detailed ablation and analysis demonstrate the importance of hierarchical register indexing, the effect and performance of view recognizer training, the good compatibility with extra complex paper search frameworks, and the online low-latency advantage compared with several powerful baselines. Code for this work is in https://github.com/Li-Z-Q/PaperRegister.
PaperRegister
As mentioned, existing paper search systems fail to handle flexible-grained queries due to they primarily collect paper abstracts to construct index of corpus, which does not contain detailed information to support queries at finer granularity. To this end, we propose PaperRegister, which supports flexible-grained paper search via hierarchical register indexing. As illustrated in Figure 2, PaperRegister consists of two stages: in the offline stage, a hierarchical index tree for paper corpus is built, and in the online stage, retrieval is performed by adaptively using suitable indexes from the tree.
Task Formulation
The paper search task involved in this work provides a query $q$ as input, with the goal of retrieving $M$ relevant papers $\{p^{(m)}\}_{m=1}^{M}$ based on the index $\mathcal{I}$ of paper corpus $\mathcal{C}$, which can be expressed as the following formula:
$\{p^{(m)}\}_{m=1}^{M}=\mathcal{F}(q,\mathcal{I})$ (1)
For regular paper search, the view of $q$ is typically a coarse-grained topic that user interests, and systems mainly construct index $\mathcal{I}$ based on paper abstracts to match against query $q$. As user’s requirements become more flexible, the view may refer to details such as module configuration and training operation. In this context, since these detailed information typically do not appear in abstract, existing paper search systems cannot perform accurate retrieval. Therefore, we propose PaperRegister to boost flexible-grained paper search, which transforms abstract-based index into hierarchical index tree for paper search, the specific process will be further explained in the following sections.
Offline Hierarchical Indexing
Considering that the granularity of view in queries can be flexible across very fine-grained, moderately fine-grained, to coarse-grained, PaperRegister offline constructs hierarchical index tree for paper corpus to support flexible paper search. Specifically, based on a hierarchical register schema, PaperRegister obtains hierarchical register for each paper through fine-grained content extracting and bottom-up content aggregating. Then the hierarchical index tree is constructed by merging hierarchical register of all papers.
Hierarchical Register Schema.
To obtain the hierarchical register for each paper, we design a hierarchical register schema, consisted of information nodes as following:
$N_{i,j}=\{n_{i,j}:(c_{i,j},~\{N_{i+1,j^{\prime}}\}_{j^{\prime}=1}^{Z})\}$ (2)
where $N_{i,j}$ is the $j$-th information node at the $i$-th layer, $n_{i,j}$ is the node name, $c_{i,j}$ is the node content that will be obtained via specific paper text, $N_{i+1,j^{\prime}}$ is a sub-node of $N_{i,j}$, $Z$ is the number of sub-nodes under each node, and $i\in[1,...,L]$, where $L$ is the number of layers in schema.
In this hierarchical register schema, the shallow-layer nodes represent coarse-grained information of the paper, and deep-layer nodes represent finer-grained information. For example, in the i+1i\!+\!1-th layer, $N_{i+1,j\prime}$ and $N_{i+1,j\prime+1}$ respectively denote module configuration and training operation in the paper, which are two kinds of relatively fine-grained information. While their common parent in the $i$-th layer, $N_{i,j}$, denotes method implementation, which is a coarser-grained summary of $N_{i+1,j\prime}$ and $N_{i+1,j\prime+1}$, thereby forming the register schema with hierarchical information nodes.
In addition, since different types of papers have different styles and formats, we design five kinds of hierarchical register schema, corresponding to five types of paper including algorithm innovation, benchmark construction, mechanism exploration, survey, and theory proof. We use a large language model to determine the type of input paper and then assign the corresponding hierarchical register schema. The detailed illustration is presented in the Appendix A.
Fine-grained Content Extracting.
Based on above hierarchical register schema, in order to obtain the fine-grained contents as much as possible, PaperRegister first perform extracting for each information node at the $L$-th layer based on paper $p^{(m)}$, which can be represented as follows:
$c_{L,j}^{(m)}=\mathcal{M}_{extract}(p^{(m)},n_{L,j})$ (3)
where $n_{L,j}$ represents the name of information node $N_{L,j}$ and $\mathcal{M}_{extract}$ is the extracting module for this process.
To achieve accurate extraction, learning from several widely recognized works, in which large language models are proven to possess reliable capabilities for content extraction (Edge et al. 2024; Li et al. 2025c), PaperRegister uses a large language model as the extracting module. Specifically, PaperRegister takes the node name and the text-formatted paper as input, uses instructions to guide in outputting corresponding content for the information node, and leaves it blank if the paper does not include corresponding content. The detailed process and instructions are in Appendix B.
Bottom-up Content Aggregating.
After extracting all the finest-grained contents, in order to obtain the complete hierarchical register, PaperRegister aggregates node contents layer by layer from bottom to top based on the hierarchical register schema, which can be represented as follows:
$c_{i,j}^{(m)}=\mathcal{M}_{aggregate}(\{c_{i+1,j\prime}^{(m)}\}_{j\prime=1}^{Z})$ (4)
where layer $i$ is from L1L-1 to 1, thereby obtaining content for each node in the hierarchical register schema for $p^{(m)}$.
To achieve accurate aggregation, as extraction process, PaperRegister uses a large language model as the aggregating module $\mathcal{M}_{aggregate}$. Specifically, PaperRegister takes contents of sub-nodes as input and uses instructions to guide the large language model in summarizing, condensing, and removing details to turn into summary text, thus obtaining the content for upper-layer information node. The detailed process and instructions are presented in the Appendix C.
At this point, PaperRegister obtains node contents at various granular levels to compose the hierarchical register for $p^{(m)}$, and then merges register of all papers in corpus $\mathcal{C}$ to construct hierarchical index tree $\mathcal{I}_{h}$, as shown in following:
$\mathcal{I}_{h}=\{\{{\rm I}_{i,j}\}_{j=1}^{Z}\}_{i=1}^{L}=\{\{\mathcal{M}_{idx}\{c_{i,j}^{(m)}\}_{m=1}^{|\mathcal{C}|}\}_{j=1}^{Z}\}_{i=1}^{L}$ (5)
where $\mathcal{M}_{idx}$ is the indexing module such as BM25 and DPR, each ${\rm I}_{i,j}$ is an index in the hierarchical index tree $\mathcal{I}_{h}$, corresponding to a kind of view for the paper corpus.
Figure 3: Illustration of view recognizer training, including SFT and GRPO via hierarchical reward, which is calculated based on the closeness level of predicted view and golden view in the hierarchical register schema.
Online Adaptive Retrieval
Based on above offline hierarchical indexing, PaperRegister perform online adaptive retrieval via suitable indexes from the hierarchical index tree. Specifically, PaperRegister first identifies views involved in the input query, determining corresponding indexes from the hierarchical index tree, then conduct retrieval by matching query against these indexes.
View Identifying.
In order to achieve adaptive retrieval, PaperRegister first uses a view recognizer to identify the view $v_{k}$ in the input query $q$, represented as follows:
$\{{v}_{k}\}_{k=1}^{K}=\mathcal{M}_{identify}(q)$ (6)
where the candidate set of ${v}_{k}$ is all node paths in the hierarchical register schema, $\mathcal{M}_{identify}$ is employed base on a small-scale language model with special training, which will be explained in next section. And to ensure that identifying results can cover the real views of query as more as possible, $\mathcal{M}_{identify}$ uses the beam search strategy (Holtzman et al. 2020) to sample the top-$K$ output views.
View-based Matching.
After identifying views in the input query, PaperRegister looks up corresponding indexes from the hierarchical index tree $\mathcal{I}_{h}$, as following:
$\{{\rm I}_{k}\}_{k=1}^{K}=\mathcal{M}_{lookup}(\mathcal{I}_{h},\{{v}_{k}\}_{k=1}^{K})$ (7)
where each ${\rm I}_{k}$ is consisted of $\{c_{k}^{(m)}\}_{m=1}^{|\mathcal{C}|}$ as in Formula 5.
Then these indexes are used to calculate relevance score of input query $q$ and each paper $p^{(m)}$, as following:
$s(q,p^{(m)})=\max\{\mathcal{M}_{rel}(q,~{c}_{k}^{(m)})_{k=1}^{K}\}$ (8)
where $\mathcal{M}_{rel}$ is relevance module such as BM25 and DPR.
Finally, the retrieved papers with top-$M$ relevance score, $\{p^{(m)}\}_{m=1}^{M}$, are selected as retrieval results for query $q$.
View Recognizer Training
For the PaperRegister system mentioned above, the view recognizer is a key aspect, which must both alleviate the latency and ensure the accuracy of view identifying. To this end, as shown in Figure 3, we use a small-scale language model with 0.6B parameters as the base model, followed by initial supervised fine-tuning (SFT), and then further enhance its view identifying capability by hierarchical-reward group relative policy optimization (GRPO) training.
Training Data.
The format of training dataset to support SFT and GRPO is shown as follows:
$\mathcal{D}_{train}=\{q_{j},v_{j}\}_{j=1}^{|\mathcal{D}_{train}|}$ (9)
where $q_{j}$ is a query, $v_{j}$ is the golden view of query, which is a node path in hierarchical register schema. To ensure the stability of training, we make the approximate same number of each view in the training dataset through manual sampling.
Supervised Fine-tuning.
To give the small-scale view recognizer basic identifying ability and make it easier for subsequent reinforcement learning training, we first perform SFT by minimizing the following loss:
$\mathcal{L}_{SFT}(\theta)=-\mathbb{E}_{(q,v)\sim\mathcal{D}_{train}}\sum_{t=1}^{|v|}\log\pi_{\theta}(v_{t}|q,v_{<t})$ (10)
where $\pi_{\theta}$ is the model and $v_{t}$ is the $t$-th token of golden view.
Hierarchical-reward GRPO.
To further strengthen the view identifying ability of small-scale view recognizer, we conduct GRPO (Shao et al. 2024) training with hierarchical reward, which is calculated based on the node path’s closeness level of predicted view and golden view in the hierarchical register schema. Specially, considering the hierarchical feature of register schema, different wrong predicted views should not be treated equally. For example, if the golden view’s path in the hierarchical register schema is Abstract-Method-Implementation-Operation, and two predicted views are Abstract-Method-Implementation-Module and Abstract-Experiment-Dataset, although both are wrong, the first one is better than the second one because it is closer to golden view in the hierarchical register schema, suggesting a higher information overlap with the golden view.
Based on this idea, we measure the path overlap between the predicted view $\hat{v}_{j}$ and the golden view $v_{j}$ in the hierarchical register schema, obtaining the hierarchical reward:
$r=\frac{\mathcal{M}_{overlap}(v_{j},\hat{v}_{j})}{|\hat{v}_{j}|}+\frac{\mathcal{M}_{overlap}(v_{j},\hat{v}_{j})}{|v_{j}|}$ (11)
where $\mathcal{M}_{overlap}$ returns the overlapping level of two inputs, $|\hat{v}_{j}|$ means the path length of predicted view in the hierarchical register schema and $|v_{j}|$ is that of golden view.
Then this hierarchical reward is used in GRPO training, which is by maximizing the following objective:
$\mathcal{J}_{GRPO}(\theta)={}\mathbb{E}_{\left[q\sim D_{train},\left\{v_{i}\right\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(v\mid q)\right]}\frac{1}{G}\sum_{i=1}^{G}\frac{1}{\left|v_{i}\right|}\sum_{t=1}^{\left|v_{i}\right|}\Biggl{\{}$
$\min\Biggl{[}\frac{\pi_{\theta}\left(v_{i,t}\mid q,v_{i,<t}\right)}{\pi_{\theta_{\text{old}}}\left(v_{i,t}\mid q,v_{i,<t}\right)}\hat{A}_{i,t},\text{clip}(\frac{\pi_{\theta}\left(v_{i,t}\mid q,v_{i,<t}\right)}{\pi_{\theta_{\text{old}}}\left(v_{i,t}\mid q,v_{i,<t}\right)},1-\epsilon,1+\epsilon)\hat{A}_{i,t}\Biggr{]}$
$-\beta\mathbb{D}_{\text{KL}}\left[\pi_{\theta}\parallel\pi_{\text{ref}}\right]\Biggr{\}},~\text{where}~\hat{A}_{i,t}=\frac{r_{i}-mean(\{r_{1},r_{2},...,r_{G}\})}{std(\{r_{1},r_{2},...,r_{G}\})}$
where $\pi_{\theta_{\text{old}}}$ is initialized by the SFT model, $G$ is the number of each group, $\epsilon$ and $\beta$ are two hyper-parameters.
Method BM25-based Matching DPR-based Matching
LitSearch F.g.Search-1 F.g.Search-2 F.g.Search-3 LitSearch F.g.Search-1 F.g.Search-2 F.g.Search-3
R@5 R@10 R@5 R@10 R@5 R@10 R@5 R@10 R@5 R@10 R@5 R@10 R@5 R@10 R@5 R@10
Direct Matching
Title 54.2 59.5 42.0 47.3 36.1 42.0 30.4 35.3 66.8 73.1 52.0 58.8 45.8 52.1 40.6 47.1
Abstract 67.1 71.3 63.7 69.0 57.4 62.9 54.2 59.7 77.7 83.2 69.4 74.3 62.1 68.1 58.2 63.1
Total Paper 64.2 68.9 79.7 82.4 81.4 83.5 84.2 86.2 74.8 81.9 72.9 78.7 68.8 73.8 65.8 71.0
Query Paraphrasing
Rewriting 61.9 69.6 58.4 64.2 54.1 59.8 50.9 56.3 76.0 83.2 67.1 72.6 60.4 65.2 55.2 60.9
HyDE 68.9 75.4 60.3 66.3 54.3 60.0 51.5 57.6 76.5 83.4 66.8 72.3 58.9 65.7 56.3 62.4
CSQE 69.0 73.8 59.1 64.4 54.2 58.6 51.3 55.9 77.9 81.7 65.8 71.1 60.2 65.5 55.2 60.9
Paper Splitting
$\text{Chunk}_{avg}$ 58.1 67.6 68.1 74.2 66.6 71.9 65.9 71.6 58.1 67.6 51.1 58.3 46.4 53.4 46.4 52.2
$\text{Chunk}_{max}$ 67.9 75.5 80.0 83.1 81.4 84.7 85.3 87.6 67.9 75.5 79.8 83.1 76.6 80.5 79.0 82.1
$\text{Paragraph}_{avg}$ 29.7 38.3 58.9 66.6 59.8 66.5 58.2 66.2 20.5 22.8 23.3 30.6 23.1 29.8 22.4 29.2
$\text{Paragraph}_{max}$ 64.3 70.9 73.8 78.8 75.4 79.9 80.8 83.9 79.5 85.0 79.2 82.8 76.5 81.4 78.8 81.8
Hierarchical Register Indexing
PaperRegister (Ours) 69.7 76.4 89.7 90.9 88.0 89.0 87.5 88.7 81.0 87.1 84.1 87.1 79.9 82.5 80.8 82.9
Table 1: Main experimental results on paper-searching tasks across various granularity, where the granularity is coarse-to-fine from LitSearch to F.g.Search-3. The table shows that PaperRegister is a powerful solution to addressing flexible-grained paper search and the advantages of PaperRegister become more pronounced at finer granular queries.
Experiments
Experimental Settings
Datasets.
We first use LitSearch (Ajith et al. 2024), mainly containing coarse-grained queries. And then we build a new dataset, Flexible-grained Search, covering queries across various granularity and including data for test, development, and training. The detailed statistical information of dataset is shown in Appendix D. And the metrics used and reported in experiments are R@5 (recall@5) and R@10 (recall@10).
(1) LitSearch (Ajith et al. 2024). By using GPT-4 to rewrite citations in papers and asking authors to write, this dataset constructs 597 paper search queries, mainly involving coarse-grained topics like “Where can I find research on evaluating consistency in generated summaries”.
(2) Flexible-grained Search. We collect 4,200 papers in machine learning field as corpus and perform the offline hierarchical indexing. To build test data across various granularity, we pick hierarchical register of 2,100 papers, employ Qwen3-32B to find original paper texts related to each content in register, then use Qwen3-32B to generate queries based on these texts, obtaining 5,644 test data including 3 parts: general-granularity (F.g.Search-1), fine-granularity (F.g.Search-2), and very fine-granularity (F.g.Search-3). The format of each test data is like $\{q_{j},p_{j}\}$. To build training and development data, we use register of another 2,100 paper and directly generate queries for contents in each paper register, where each query is with a golden view, the format of training and development data is like $\{q_{j},v_{j},p_{j}\}$. Then we do random splitting and get 13,824 training data and 695 development data, the training data are used in view recognizer training part as illustrated in the Formula 9 and development data are used for analysis in the Table 3.
Selected Baselines.
We select baselines from commonly used or recent methods suitable for paper search tasks.
(1) Direct Matching. These methods directly use paper titles, abstracts or total text of papers to construct index of corpus, and retrieval by matching query with these contents.
(2) Query-paraphrasing Methods. These methods aim to improve paper‐searching performance by paraphrasing the input query, including: Rewriting (Ma et al. 2023), which uses a LLM to rewrite original query. HyDE (Gao et al. 2023), which uses a LLM to generate a fake document based on input query and retrieves real documents by this fake document. CSQE (Lei et al. 2024), which initially retrieves several documents and then uses a LLM to expand original query based on these initially-retrieved documents. We use Qwen3-32B as the LLM when conducting these baselines.
(3) Paper-splitting Methods. These methods split the original paper into multiple short parts, then calculate the similarity between the query and each short part, and finally integrate all the similarity to determine the overall similarity of query and original paper. In experiments, we employ four kinds of settings, including splitting the original paper to the fixed length of 512-token chunk or by the raw paragraph, taking the average similarity or the maximum similarity of all parts as the final score, represented as $\text{Chunk}_{avg}$, $\text{Chunk}_{max}$, $\text{Paragraph}_{avg}$, and $\text{Paragraph}_{max}$, respectively.
Method BM25-based Matching DPR-based Matching
LitSearch F.g.Search-1 F.g.Search-2 F.g.Search-3 LitSearch F.g.Search-1 F.g.Search-2 F.g.Search-3
R@5 R@10 R@5 R@10 R@5 R@10 R@5 R@10 R@5 R@10 R@5 R@10 R@5 R@10 R@5 R@10
PaperRegister (Ours) 69.7 76.4 89.7 90.9 88.0 89.0 87.5 88.7 81.0 87.1 84.1 87.1 79.9 82.5 80.8 82.9
w/ only layer-1 64.5 70.4 87.8 90.6 77.4 80.2 73.5 76.7 79.0 83.5 82.5 85.2 73.4 77.1 68.1 72.2
w/ only layer-2 66.8 73.8 84.4 86.5 87.1 88.8 82.0 84.4 78.9 84.2 81.1 84.9 79.3 82.2 73.5 77.9
w/ only layer-3 64.6 71.7 79.9 81.8 83.7 85.0 86.8 88.1 78.8 85.2 77.8 81.1 77.3 80.0 80.4 82.6
Table 2: Ablation results for the hierarchical register indexing. The table shows that hierarchical index tree are important for flexible-grained paper search and different layers in the hierarchical register schema serve queries at different granularity.
Implementation Details.
In the offline stage, we employ Qwen3-32B (Qwen 2025) as the large language model in process of fine-grained content extracting and bottom-up content aggregating, and deploy it as API by vllm [^1] https://pypi.org/project/vllm/ on two A-100 80G GPU for convenience. In the online stage, we set $K$ as 5 and $M$ as 5 or 10, use Qwen3-0.6B (Qwen 2025) as the base model of view recognizer, employ a prefix tree-based restricted decoding strategy (Tang et al. 2024) in the view identifying process to prevent the model outputting irrelevant token, and conduct the view-based matching process via rank-bm25222https://pypi.org/project/rank-bm25/ for BM25 and gte-Qwen2-7B-instruct (Li et al. 2023) for DPR. For the view recognizer training, we use TRL [^3] https://github.com/huggingface/trl framework to conduct SFT and GRPO. We set train epoch as 5 in SFT, $G$ as 5 and train epoch as 2 in GRPO, and keep all other parameters as default value in TRL. When conducting baselines, we keep the same settings as above.
Overall Results
Results compared with baselines are shown in the Table 1, there are two main conclusions based on the results:
(1) PaperRegister is a powerful solution to addressing flexible-grained paper search tasks. As shown in the Table 1, PaperRegister demonstrates excellent performance in paper search tasks at various granularity, outperforming all baselines in both BM25-based matching and DPR-based matching settings, on both recall@5 and recall@10 metrics. For example, in F.g.Search-3, under the DPR-based matching setting, PaperRegister achieves a recall@5 score of 80.8, while using abstract-based index yields only 58.2, where PaperRegister can achieve a performance improvement of 22.6. All in all, PaperRegister can improve a lot compared with various previous paper search methods.
Figure 4: Paper search performance of the PaperRegister with different view recognizer. The figure shows strong recognizer is with obvious positive impact on overall system.
(2) Compared to traditional methods, the advantages of PaperRegister become more pronounced at these finer granular queries. According to experimental setting, from LitSearch to F.g.Search-1 to F.g.Search-2 to F.g.Search-3, the granularity of the paper search query becomes increasingly finer. And the experimental results show that the performance improvement of PaperRegister over the method using abstract-based index becomes larger. For example, under the DPR-based matching setting, PaperRegister achieves recall@5 scores of 81.0, 84.1, 79.9, and 80.8 on the four datasets, respectively, while using the abstract-based index yields scores of 77.7, 69.4, 62.1, and 58.2. Here, PaperRegister delivers performance improvements of 3.3, 14.7, 17.8, and 22.6, respectively. Therefore, PaperRegister exhibits more improvements over baseline on finer-grained queries, which can to some extent validate rationality of this work’s motivation, that traditional paper search via abstract-based index struggles to handle fine-grained queries.
Ablation Results
To validate the importance of hierarchical register indexing, we employ ablation experiments by simplifying the hierarchical register schema. As shown in Table 2, where “w/ only layer-1” means retaining only the coarsest-grained nodes, “w/ only layer-2” means retaining only the medium-grained nodes, and “w/ only layer-3” means retaining only the finest-grained nodes. Based on these simplified register schema for experiments, we get the following conclusions:
(1) Hierarchical register indexing is important for ensuring the effectiveness of paper search. As shown in Table 2, compared to using the complete hierarchical register schema, using a register schema composed of nodes from any single layer leads to obvious performance drop. For example, recall@5 of LitSearch under BM25-based matching drops from 69.7 to 64.5, 66.8, 64.6 for “w/ only layer-1”, “w/ only layer-2”, “w/ only layer-3”, respectively. And other datasets also show similar situation. Therefore, results strongly prove importance of hierarchical register indexing.
(2) Different layers in the hierarchical register schema serve queries at different granularity. For example, under “w/ only layer-1” setting, the DPR-based matching and recall@5 metric of F.g.Search-1 drops from 84.1 to 82.5 (1.6 decrease), while F.g.Search-1 is from 80.8 to 68.1 (12.7 decrease), which is significantly larger than 1.6. Conversely, under “w/ only layer-3” setting, F.g.Search-1 drops from 84.1 to 77.8 (6.3 decrease), while F.g.Search-3 is from 80.8 to 80.4 (0.4 decrease), which is much smaller than 6.3. Therefore, nodes from different layers play specific roles in handling queries of different granularity, further strongly demonstrating necessity of hierarchical register indexing.
Recognizer ACC $\uparrow$ Latency (s) $\downarrow$
Qwen3-32B 30.5 28.3
Qwen3-32B (few-shot) 47.8 37.8
View Recognizer in PaperRegister 83.5 2.3
w/   only SFT 80.9 -
w/o hierarchical reward 81.7 -
Table 3: Comparison and ablation for the view recognizer training. The table shows that the training process in this work is an effective approach to obtain a view recognizer with both high performance and low latency.
Detailed Analysis
In this section, we conduct detailed analysis including: the effect of view recognizer, the compatibility of PaperRegister with complex PaSa framework (He et al. 2025), and an analysis of the online search efficiency. Due to the space constraint, we only report the performance that is based on DPR-based matching for analysis in this section.
Effect of View Recognizer and Training.
To analyze the role of view recognizer in PaperRegister and necessity of our training, we first examine relationship between recognizer capability and PaperRegister performance in Figure 4, then compare with Qwen3-32B and conduct ablation in Table 3.
As shown in Figure 4, where “Bad Recognizer”,“Random Recognizer”,“Weak Recognizer” represent the completely incorrect, randomly predicting, and weak-performing view recognizers, respectively, the curve demonstrates a clear positive correlation between the capability of view recognizer and final performance of PaperRegister. Therefore, building an accurate view recognizer is a key factor for the PaperRegister to achieve high-performance paper search.
As shown in Table 3, We compare accuracy and latency of view identifying between the view recognizer in PaperRegister and Qwen3-32B (enable thinking) under zero-shot and few-shot settings. The results show the view recognizer in PaperRegister is with better accuracy and latency than powerful Qwen3-32B. Furthermore, we only keep SFT for the view recognizer training, or replace the hierarchical reward with direct 0-1 reward in GRPO training. Results show that the accuracy decreases to some extent. Therefore, training a small-scale language model with SFT and then hierarchical-reward GRPO is an effective approach to obtain the view recognizer with both high performance and low latency.
Compatibility with Complex PaSa Framework.
Considering that PaperRegister is mainly a retrieval module, and given that some recent complex paper search frameworks such as PaSa (He et al. 2025) incorporate various modules like rewriting, retrieval, iteration, and filtering, in order to explore whether PaperRegister can be compatible with such complex frameworks and further enhance performance, we replace the original retrieval module in PaSa with PaperRegister. As shown in Figure 5, PaperRegister can further improve the performance of PaSa across paper search tasks at various granularity. Therefore, PaperRegister can be effectively adapted as a retrieval module into complex paper search frameworks and further enhance the capability.
Figure 5: Performance of adding PaperRegister into the PaSa framework. The figure shows that PaperRegister can greatly cooperate with extra complex modules in PaSa framework.
Analysis for Online Search Efficiency.
Considering that the efficiency of online paper search system is a crucial aspect of its usability in practical scenarios, we compare the online search efficiency of PaperRegister with multiple baselines. As shown in the Table 4, the online latency of these baselines is significantly higher than that of PaperRegister, which limits the practical applicability of these methods in real-world scenarios. In contrast, PaperRegister demonstrates the acceptable online latency, proving its good potential for application in the real-world paper search.
Related Work
To achieve accurate paper search, previous related works mainly focus on enhancing the input query, with less attention given to the index of paper corpus. Some works improve paper search performance by rewriting the original query, such as using powerful large language models for rewriting based on some manually designed instructions (Ma et al. 2023; Anand et al. 2023), generating pseudo-documents to replace original query for retrieval real documents (Gao et al. 2023; Li et al. 2024a), or employing several powerful models as agent system to conduct multi-round rewriting and expansion (He et al. 2025; Ren et al. 2025). Another kind of works expands the input query through pseudo-relevance feedback, such as augmenting the query based on initially retrieved documents (Lei et al. 2024; Li et al. 2025a), or extracting several keywords from the corpus to enrich the information of original query (Kang et al. 2024; Zhang et al. 2025). Although these works can improve performance in regular paper search, they fail to address flexible-grained queries because they do not solve the flaws in the index of paper corpus. Recently, several works split original paper by fine-grained fields and then construct multiple-field indexes for retrieval (Sotaro et al. 2024; Li et al. 2025b; Shi et al. 2025; Chen et al. 2025). However, these works still cannot effectively handle flexible-grained queries because they only focus on a single granularity level.
Compared with above related works, the largest advantage of PaperRegister is hierarchical register indexing and conducting paper search with hierarchical index tree, which can effectively handle flexible-grained paper search.
Method Online Paper Search Latency (s) $\downarrow$
Rewriting 9.3
HyDE 20.7
CSQE 33.5
$\text{Chunk}_{max}$ 5.4
$\text{Paragraph}_{max}$ 4.2
PaperRegister (Ours) 2.5
Table 4: Comparison of online latency among PaperRegister and several baselines. The table show that PaperRegister is with better real-world applicability with less latency.
Conclusion
In this work, noticed limitation of existing methods in paper search tasks, we propose PaperRegister, boosting flexible-grained paper search via hierarchical register indexing. In the offline stage, PaperRegister constructs a hierarchical index tree by fine-grained content extracting and bottom-up content aggregating. In the online stage, PaperRegister first identify views of input query, then matches the query against corresponding indexes from the hierarchical index tree. Furthermore, we construct the view recognizer by applying supervised finetuning and hierarchical-reward group relative policy optimization to a small-scale language model. Experiments on extensive flexible-grained paper search tasks demonstrate that PaperRegister is an effective solution, which reaches the SOTA performance, with improvement being more pronounced as query granularity finer. Therefore, this work offers a promising direction for developing more powerful paper search systems in future.
Appendix A
This part is about detailed illustration for the hierarchical register schema.
Figure 6: The first kind of hierarchical register schema.
Figure 7: The second kind of hierarchical register schema.
Figure 8: The third kind of hierarchical register schema.
Figure 9: The fourth kind of hierarchical register schema.
Figure 10: The fifth kind of hierarchical register schema.
Since different types of papers have the different styles and formats, we design five kinds of hierarchical register schema, corresponding to five types of paper including algorithm innovation, benchmark construction, mechanism exploration, survey, and theory proof. We use a large language model to determine the type of input paper and then assign the corresponding hierarchical register schema. The details of each kind of schema is in Figure 6 7 8 9 10. And the instruction used to help LLM to do determine is ==Instruction== Determine the category of the paper based on its abstract. The category options are as follows: Algorithm Innovation: Proposes new systems, new models, new training methods, new inference approaches, new data organization methods, etc. Benchmark Construction: Introduces a new benchmark. Mechanism Exploration: Investigates and analyzes the mechanisms of existing systems, algorithms, phenomena, or functionalities. Theory Proof: Proves a certain theory or formula. Survey and Review: Summarizes the research landscape of a particular field. Your final output should be only the best correct category of the paper, do not contain any other information or explanation. ==Abstract== abstract
Appendix B
This part is about detailed process and instructions for the fine-grained content extracting.
Learning from several widely recognized works, in which large language models are proven to possess reliable capabilities for content extraction (Edge et al. 2024; Li et al. 2025c, 2024b, 2024c), PaperRegister uses a large language model as the extracting module. PaperRegister takes the node name and the text-formatted paper as input, uses instructions to guide in outputting corresponding content for the information node, and leaves it blank if the paper does not include corresponding content. In addition, PaperRegister also retrieve relevant paper text after the extracting as supplement to improve the register content. The instruction used for LLM is ==Instruction== You are an content extraction expert, particularly skilled at extracting structured records from academic papers. Below, I need you to perform extraction based on a given schema. Please note the following: 1. Your extraction does not need to preserve the original text verbatim. You may paraphrase or summarize the content to make the extracted content more comprehensive and fluent. 2. Not all field names in the schema will have corresponding content in the paper. If you cannot find a precise match in the paper, leave that field empty. 3. Your final output must strictly adhere to the original schema in JSON format, starting with ’“‘json’ and ending with ’“‘’. ==Schema== schema ==Paper== paper
Appendix C
This part is about detailed process and instructions for the bottom-up content extracting.
PaperRegister takes contents of sub-nodes as input and uses instructions to guide the large language model in summarizing, condensing, and removing details to turn into summary text, thus obtaining the content for upper-layer information node. In addition, PaperRegister also retrieve relevant paper text after the extracting as supplement to improve the register content. The instruction used for LLM is ==Instruction== You are an information integration expert, and now I need your help to complete an information integration task. I will provide you with a two-level tree structure, including a root node and two child nodes. Ideally, the content of the root node should be a summary and generalization of the two child nodes. Your task is to generate the content of the root node based on the content of the two child nodes. Note the following: 1. The input I provide you is a dictionary in JSON format, including the keys ‘root_name‘ and ‘children‘, where the value of ‘children‘ is a list of child nodes. 2. Each child node contains three fields: ‘node_name‘, ‘node_desc‘, and ‘node_value‘. You should primarily use the content of ‘node_value‘ for summarization and generalization. 3. The ‘root_value‘ field should provide an abstraction and summary of the two child nodes’ contents. In other words, the root_value must not repeat keywords from the child nodes; instead, it should abstract based on those keywords. More strictly, the root_value length must not exceed that of either child node. 4. Your output should be a dictionary in JSON format, meaning the input dictionary will have content for the ‘root_value‘ field. The final output should start with ’“‘json’ and end with ’“‘’. ==Input Tree== tree
Appendix D
This part is about detailed statistical information of dataset for the experiments.
We first use LitSearch, mainly containing coarse-grained queries. And then we build a new dataset, Flexible-grained Search, covering queries across various granularity and including data for test, development, and training. Due to limited computational resources, we are unable to conduct experiments based on an ultra-large-scale paper corpus. For LitSearch, we extract 3,400 papers as the corpus, using the original 597 paper search queries. For Flexible-grained Search, we collect 4,200 papers as the corpus. The numbers of queries included in F.g.Search-1, F.g.Search-2, and F.g.Search-3 are 1,922, 1,922, and 1,800, respectively.

### v2

@@
 Abstract
-Paper search is an important activity for researchers, typically involving using a query with description of a topic to find relevant papers.
-As research deepens, paper search requirements may become more flexible, sometimes involving specific details such as module configuration rather than being limited to coarse-grained topics.
-However, previous paper search systems are unable to meet these flexible-grained requirements, as these systems mainly collect paper abstracts to construct index of corpus, which lack detailed information to support retrieval by finer-grained queries.
-In this work, we propose PaperRegister, consisted of offline hierarchical indexing and online adaptive retrieval, transforming traditional abstract-based index into hierarchical index tree for paper search, thereby supporting queries at flexible granularity.
-Experiments on paper search tasks across a range of granularity demonstrate that PaperRegister achieves the state-of-the-art performance, and particularly excels in fine-grained scenarios, highlighting the good potential as an effective solution for flexible-grained paper search in real-world applications.
-Introduction
-Paper search is an important and almost everyday activity for researchers (Kuhlthau 1991; Ellis, Cox, and Hall 1993; Hemminger et al.
-2007; Case and Given 2016).
-Typically, this process begins when the user submits a natural-language query describing a topic, then the retrieval system matches this query against the paper corpus and returns the subset of papers with the highest relevance (Wadden et al.
-2020; Cohan et al.
-2020; Ajith et al.
-2024; He et al.
-2025).
-As researchers delve more deeply into their work, the paper search requirements can become increasingly flexible.
-For example, the view of input query may refer to detailed module configuration or methodological operation, rather than being limited to the level of coarse-grained topic (Mysore et al.
-2021; Kang et al.
-2024; Wang et al.
-2023; Zhang et al.
-2025).
-Therefore, a paper search system supporting queries at flexible levels of granularity is of important value.
-Unfortunately, existing paper search systems cannot effectively handle queries across flexible granularity, since they mainly use paper abstracts to construct index of corpus for retrieval (Zheng et al.
-2020; Gao et al.
-2023; Mackie, Chatterjee, and Dalton 2023; Lei et al.
-2024).
-When a query’s view involves finer‐grained information that does not appear in abstract, they fails to retrieve relevant papers.
-As shown in Figure 1, the input query “Retrieve papers about jointly training encoder and generator via minimizing the negative marginal log-likelihood without document encoder” specifies a detailed training operation, expecting to find the relevant paper.
-However, the target cannot be successfully retrieved by traditional paper search systems via abstract-based index, because the abstract does not mention such detailed training operation.
-Therefore, building a powerful paper search system that supports queries at flexible granularity remains an important open challenge.
-Figure 1: PaperRegister can support flexible-grained paper search through the hierarchical register, while traditional method fails due to abstract cannot contain required details.
-Figure 2: The overview of PaperRegister, including hierarchical indexing and adaptive retrieval.
-In offline stage, PaperRegister constructs hierarchical index tree via fine-grained content extracting and bottom-up content aggregating based on hierarchical register schema.
-In online stage, PaperRegister first identify the views of input query and then conduct view-based matching.
-To this end, we propose PaperRegister, which transforms traditional abstract-based index into hierarchical index tree for paper search, thereby supporting queries at flexible granularity.
+As researchers delve more deeply into their work, paper search requirements may become more flexible, sometimes involving specific details such as module configuration rather than being limited to coarse-grained topics.
+However, previous paper search systems are unable to meet these flexible-grained requirements, as previous systems mainly collect paper abstract to construct corpus index, which lacks detailed information to support retrieval by some finer-grained queries.
+In this work, we propose PaperRegister, which transforms traditional abstract-based index into a hierarchical index tree, thereby supporting queries at flexible granularity.
+Experiments on paper search tasks across a range of granularity demonstrate that PaperRegister achieves the SOTA performance, and particularly excels in the fine-grained scenarios, highlighting good potential as an effective solution for flexible-grained paper search in real-world applications.
+https://github.com/Li-Z-Q/PaperRegister.
+PaperRegister: Boosting Flexible-grained Paper Search via
+Hierarchical Register Indexing
+Zhuoqun Li1,2, Xuanang Chen1, Hongyu Lin1, Yaojie Lu1, Xianpei Han1, Shanshan Jiang3, Bin Dong3, Le Sun1 1Chinese Information Processing Laboratory, Institute of Software, Chinese Academy of Sciences 2University of Chinese Academy of Sciences 3Ricoh Software Research Center Beijing, Ricoh Company, Ltd.
+{lizhuoqun2021,chenxuanang,hongyu,luyaojie,xianpei,sunle}@iscas.ac.cn
+1 Introduction
+Paper search is an important and almost everyday activity for researchers kuhlthau1991inside; ellis1993comparison; hemminger2007information; case2016looking.
+Typically, this process begins when user submits a natural-language query describing a topic, and then retrieval system matches this query against paper corpus and returns subset of papers with the highest relevance wadden2020fact; cohan2020specter; ajith2024litsearch; he2025pasa.
+As researchers delve more deeply into their work, paper search requirements can become increasingly flexible.
+For example, the view of query may refer to detailed module configuration or methodological operation, rather than being limited to the level of coarse-grained topic mysore2021csfcube; kang2024taxonomy; wang2023scientific; zhang2025scientific.
+Therefore, a paper search system supporting queries at flexible granularity is of important value.
+Figure 1: PaperRegister supports flexible-grained paper search via hierarchical register, while traditional method fails due to abstract cannot contain required details.
+Unfortunately, existing paper search systems cannot effectively handle queries across flexible granularity, since they mainly use paper abstract to construct corpus index for retrieval zheng2020bert; gao2023precise; mackie2023generative; lei2024corpus.
+When the view of query involves finer‐grained information that does not appear in abstract, they fail to retrieve relevant papers.
+As shown in Figure 1, the query “Retrieve papers about jointly training encoder and generator via minimizing negative marginal log-likelihood without doc encoder” focuses on a detailed training operation.
+For this query, the target papers cannot be successfully retrieved by traditional abstract-based index, because paper abstract does not mention such a detailed training operation.
+Therefore, building a paper search system supporting queries at flexible granularity remains a valuable and unresolved challenge.
+To this end, we propose PaperRegister, which transforms traditional abstract-based index into a hierarchical index tree, thereby supporting queries at flexible granularity.
 Specially, PaperRegister includes offline hierarchical indexing and online adaptive retrieval.
-In the offline stage, to construct a hierarchical index tree for corpus, PaperRegister uses a hierarchical register schema, consisted of information nodes at various granularity, in which each node path represents a kind of view.
-Based on this schema, PaperRegister first extracts fine‐grained contents for each paper, then aggregates contents bottom‐up, layer by layer, yielding hierarchical register of the paper.
-And then hierarchical index tree is constructed by merging the register of all papers.
-In the online stage, to adaptively retrieve via the hierarchical index tree, PaperRegister first employs a view recognizer to identify views of input query, which associates to suitable indexes from the hierarchical index tree, then matches query against these indexes to achieve precise paper search.
-One key aspect of PaperRegister is the view recognizer, which must be both low latency and high accuracy.
+In the offline stage, to construct the hierarchical index tree for paper corpus, PaperRegister uses a hierarchical register schema, which is consisted of information nodes at various granularity, and each node path represents a kind of view.
+Based on this schema, PaperRegister first extracts fine‐grained node contents, then aggregates node contents bottom‐up, layer by layer, yielding hierarchical register for each paper.
+And then the hierarchical index tree is constructed by merging register of all papers.
+In the online stage, to adaptively retrieve via the hierarchical index tree, PaperRegister first employs a view recognizer to identify the views of query, which associates to suitable indexes from the tree, and then matches query against these indexes to achieve the precise paper search.
+One key aspect of PaperRegister is the view recognizer, which must be both low-latency and high-accuracy.
 To this end, we construct the view recognizer by training a small-scale language model via hierarchical-reward reinforcement learning.
-Firstly, in order to alleviate the online latency, we adopt a language model with 0.6 billion parameter as base.
-Then, in order to equip the model with basic identifying capability, we perform supervised fine-tuning (SFT) on it, with query as input and the golden view as label.
-Furthermore, in order to further enhance accuracy of the view recognizer, considering the hierarchical dependency of views in the hierarchical register schema, we design a hierarchical reward as the reinforcement learning signal, which is calculated by closeness level between the predicted view and golden view in hierarchical register schema.
-And then we employ this reward on group relative policy optimization (GRPO)  (Shao et al.
-2024) training to enhance the view recognizer.
-In experiments, we compare PaperRegister with some common-used baselines on paper search tasks across various levels of granularity.
-The results show PaperRegister achieves state-of-the-art performance, with improvements being more pronounced as query granularity finer, confirming that PaperRegister is a robust solution for challenging flexible-grained paper search tasks.
-Additionally, detailed ablation and analysis demonstrate the importance of hierarchical register indexing, the effect and performance of view recognizer training, the good compatibility with extra complex paper search frameworks, and the online low-latency advantage compared with several powerful baselines.
-Code for this work is in https://github.com/Li-Z-Q/PaperRegister.
-PaperRegister
-As mentioned, existing paper search systems fail to handle flexible-grained queries due to they primarily collect paper abstracts to construct index of corpus, which does not contain detailed information to support queries at finer granularity.
-To this end, we propose PaperRegister, which supports flexible-grained paper search via hierarchical register indexing.
-As illustrated in Figure 2, PaperRegister consists of two stages: in the offline stage, a hierarchical index tree for paper corpus is built, and in the online stage, retrieval is performed by adaptively using suitable indexes from the tree.
-Task Formulation
-The paper search task involved in this work provides a query $q$ as input, with the goal of retrieving $M$ relevant papers $\{p^{(m)}\}_{m=1}^{M}$ based on the index $\mathcal{I}$ of paper corpus $\mathcal{C}$, which can be expressed as the following formula:
+Firstly, to alleviate online latency, we adopt a language model with 0.6 billion parameters as base.
+Then, to equip the model with basic identifying ability, we perform supervised fine-tuning, where query is the input and golden view is the label.
+Furthermore, to enhance accuracy of view recognizer, considering the hierarchical dependency of nodes in register schema, we design a hierarchical reward as reinforcement learning signal, which is calculated by the closeness level between predicted view and golden view in hierarchical register schema.
+And then we employ this reward on group relative policy optimization (GRPO)  Shao2024DeepSeekMathPT to enhance the capability of view recognizer.
+Figure 2: PaperRegister includes hierarchical indexing and adaptive retrieval.
+Offline, PaperRegister constructs hierarchical index tree via fine-grained content extracting and bottom-up content aggregating based on a hierarchical register schema.
+Online, PaperRegister first identify views of query and then conduct view-based matching.
+In experiments, we compare PaperRegister with several common-used methods on paper search, across various levels of granularity.
+Results show PaperRegister achieves the SOTA performance, with the improvement being more pronounced as query granularity is finer, confirming that PaperRegister is a robust solution for the challenging flexible-grained paper search task.
+The main contributions of this work can be summarized as follows:
+  • We construct PaperRegister, which can support paper search queries at flexible granularity via the hierarchical register indexing.
+  • We design hierarchical-reward reinforcement learning, which can train a powerful and low-latency view recognizer for PaperRegister.
+  • We conduct extensive experiments and analysis, which show PaperRegister is an advanced system for flexible-grained paper search.
+2 PaperRegister
+Existing paper search systems fail to handle queries at flexible granularity due to they primarily collect paper abstract to construct index of corpus, which does not contain detailed information to support queries at finer granularity.
+To this end, we propose PaperRegister, which can support flexible-grained paper search via hierarchical register indexing.
+As illustrated in Figure 2, in the offline stage, a hierarchical index tree is built for paper corpus, and in the online stage, adaptive retrieval is performed by selecting and using suitable indexes from the tree.
+2.1 Task Formulation
+The task involved in this work provides a query $q$ as input, with the goal of retrieving $M$ relevant papers $\{p^{(m)}\}_{m=1}^{M}$ based on the index $\mathcal{I}$ of paper corpus $\mathcal{C}$, which can be expressed as follows:
 $\{p^{(m)}\}_{m=1}^{M}=\mathcal{F}(q,\mathcal{I})$ (1)
-For regular paper search, the view of $q$ is typically a coarse-grained topic that user interests, and systems mainly construct index $\mathcal{I}$ based on paper abstracts to match against query $q$.
+For regular paper search, view of query $q$ is typically a coarse-grained topic that user is interested in, and systems mainly construct index $\mathcal{I}$ based on paper abstract to match against $q$.
 As user’s requirements become more flexible, the view may refer to details such as module configuration and training operation.
-In this context, since these detailed information typically do not appear in abstract, existing paper search systems cannot perform accurate retrieval.
-Therefore, we propose PaperRegister to boost flexible-grained paper search, which transforms abstract-based index into hierarchical index tree for paper search, the specific process will be further explained in the following sections.
-Offline Hierarchical Indexing
-Considering that the granularity of view in queries can be flexible across very fine-grained, moderately fine-grained, to coarse-grained, PaperRegister offline constructs hierarchical index tree for paper corpus to support flexible paper search.
-Specifically, based on a hierarchical register schema, PaperRegister obtains hierarchical register for each paper through fine-grained content extracting and bottom-up content aggregating.
-Then the hierarchical index tree is constructed by merging hierarchical register of all papers.
+In this context, since this detailed information typically do not appear in abstract, existing paper search systems cannot retrieve accurately.
+2.2 Offline Hierarchical Indexing
+Considering the granularity of query view can be flexible across very fine-grained, moderately fine-grained, to coarse-grained, PaperRegister offline constructs a hierarchical index tree for paper corpus to support flexible paper search.
+Specifically, based on a hierarchical register schema, PaperRegister obtains hierarchical register for each paper through fine-grained content extracting and bottom-up content aggregating, and then hierarchical index tree is constructed by merging register of all papers.
 Hierarchical Register Schema.
-To obtain the hierarchical register for each paper, we design a hierarchical register schema, consisted of information nodes as following:
+To obtain the hierarchical register for each paper, we first design a hierarchical register schema, which consists of information nodes as the following formula:
 $N_{i,j}=\{n_{i,j}:(c_{i,j},~\{N_{i+1,j^{\prime}}\}_{j^{\prime}=1}^{Z})\}$ (2)
-where $N_{i,j}$ is the $j$-th information node at the $i$-th layer, $n_{i,j}$ is the node name, $c_{i,j}$ is the node content that will be obtained via specific paper text, $N_{i+1,j^{\prime}}$ is a sub-node of $N_{i,j}$, $Z$ is the number of sub-nodes under each node, and $i\in[1,...,L]$, where $L$ is the number of layers in schema.
-In this hierarchical register schema, the shallow-layer nodes represent coarse-grained information of the paper, and deep-layer nodes represent finer-grained information.
-For example, in the i+1i\!+\!1-th layer, $N_{i+1,j\prime}$ and $N_{i+1,j\prime+1}$ respectively denote module configuration and training operation in the paper, which are two kinds of relatively fine-grained information.
-While their common parent in the $i$-th layer, $N_{i,j}$, denotes method implementation, which is a coarser-grained summary of $N_{i+1,j\prime}$ and $N_{i+1,j\prime+1}$, thereby forming the register schema with hierarchical information nodes.
-In addition, since different types of papers have different styles and formats, we design five kinds of hierarchical register schema, corresponding to five types of paper including algorithm innovation, benchmark construction, mechanism exploration, survey, and theory proof.
-We use a large language model to determine the type of input paper and then assign the corresponding hierarchical register schema.
-The detailed illustration is presented in the Appendix A.
+where $N_{i,j}$ is the $j$-th information node at the $i$-th layer, $n_{i,j}$ is node name, $c_{i,j}$ is expected node content from specific paper, $N_{i+1,j^{\prime}}$ is a sub-node of $N_{i,j}$, $Z$ is number of sub-nodes under each node, and $i\in[1,...,L]$, $L$ is number of layers in schema.
+In hierarchical register schema, shallow-layer nodes represent coarse-grained information, and deep-layer nodes represent finer-grained information.
+For example, in the i+1i\!+\!1-th layer, $N_{i+1,j\prime}$ and $N_{i+1,j\prime+1}$ respectively denote module configuration and training operation of paper, which are two kinds of relatively fine-grained information.
+While their common parent in the $i$-th layer, $N_{i,j}$, denotes method implementation, which is a coarser-grained summary of $N_{i+1,j\prime}$ and $N_{i+1,j\prime+1}$, thereby forming schema with hierarchical-dependency nodes.
+In addition, since different types of papers have different styles and formats, we design five kinds of schema, corresponding to five paper types, and use a large language model to determine the type of each paper.
+Details are presented in Appendix A.
 Fine-grained Content Extracting.
-Based on above hierarchical register schema, in order to obtain the fine-grained contents as much as possible, PaperRegister first perform extracting for each information node at the $L$-th layer based on paper $p^{(m)}$, which can be represented as follows:
+Based on the above hierarchical register schema, in order to obtain fine-grained contents as much as possible, PaperRegister first extracts content for each information node at the $L$-th layer from paper $p^{(m)}$, which can be represented as the following formula:
 $c_{L,j}^{(m)}=\mathcal{M}_{extract}(p^{(m)},n_{L,j})$ (3)
-where $n_{L,j}$ represents the name of information node $N_{L,j}$ and $\mathcal{M}_{extract}$ is the extracting module for this process.
-To achieve accurate extraction, learning from several widely recognized works, in which large language models are proven to possess reliable capabilities for content extraction (Edge et al.
-2024; Li et al.
-2025c), PaperRegister uses a large language model as the extracting module.
-Specifically, PaperRegister takes the node name and the text-formatted paper as input, uses instructions to guide in outputting corresponding content for the information node, and leaves it blank if the paper does not include corresponding content.
-The detailed process and instructions are in Appendix B.
+where $n_{L,j}$ represents name of information node $N_{L,j}$ and $\mathcal{M}_{extract}$ is the extracting module.
+To achieve accurate extraction, learning from several widely recognized works, in which large language models are proven to possess reliable capabilities for content extraction tasks edge2024local; li2025deepsolution, PaperRegister employs a large language model as extracting module.
+Specifically, PaperRegister takes the node name and text-formatted paper as input, uses instructions to guide in outputting corresponding content for the information node, and leaves it blank if the paper does not include corresponding content.
+The detailed process and instructions are in the Appendix B.
+Figure 3: Illustration of view recognizer training, including SFT and GRPO via hierarchical reward, which is calculated based on the closeness level of predicted view and golden view in the hierarchical register schema.
 Bottom-up Content Aggregating.
-After extracting all the finest-grained contents, in order to obtain the complete hierarchical register, PaperRegister aggregates node contents layer by layer from bottom to top based on the hierarchical register schema, which can be represented as follows:
+After extracting all the finest-grained contents, in order to smoothly obtain a complete hierarchical register, PaperRegister aggregates node contents layer by layer from bottom to top based on the hierarchical register schema, represented as following formula:
 $c_{i,j}^{(m)}=\mathcal{M}_{aggregate}(\{c_{i+1,j\prime}^{(m)}\}_{j\prime=1}^{Z})$ (4)
-where layer $i$ is from L1L-1 to 1, thereby obtaining content for each node in the hierarchical register schema for $p^{(m)}$.
-To achieve accurate aggregation, as extraction process, PaperRegister uses a large language model as the aggregating module $\mathcal{M}_{aggregate}$.
-Specifically, PaperRegister takes contents of sub-nodes as input and uses instructions to guide the large language model in summarizing, condensing, and removing details to turn into summary text, thus obtaining the content for upper-layer information node.
-The detailed process and instructions are presented in the Appendix C.
-At this point, PaperRegister obtains node contents at various granular levels to compose the hierarchical register for $p^{(m)}$, and then merges register of all papers in corpus $\mathcal{C}$ to construct hierarchical index tree $\mathcal{I}_{h}$, as shown in following:
-$\mathcal{I}_{h}=\{\{{\rm I}_{i,j}\}_{j=1}^{Z}\}_{i=1}^{L}=\{\{\mathcal{M}_{idx}\{c_{i,j}^{(m)}\}_{m=1}^{|\mathcal{C}|}\}_{j=1}^{Z}\}_{i=1}^{L}$ (5)
-where $\mathcal{M}_{idx}$ is the indexing module such as BM25 and DPR, each ${\rm I}_{i,j}$ is an index in the hierarchical index tree $\mathcal{I}_{h}$, corresponding to a kind of view for the paper corpus.
-Figure 3: Illustration of view recognizer training, including SFT and GRPO via hierarchical reward, which is calculated based on the closeness level of predicted view and golden view in the hierarchical register schema.
-Online Adaptive Retrieval
-Based on above offline hierarchical indexing, PaperRegister perform online adaptive retrieval via suitable indexes from the hierarchical index tree.
-Specifically, PaperRegister first identifies views involved in the input query, determining corresponding indexes from the hierarchical index tree, then conduct retrieval by matching query against these indexes.
+where the layer $i$ is from L1L-1 to 1, thereby obtaining content for each information node in the hierarchical register schema for the paper $p^{(m)}$.
+To achieve accurate aggregation, like extraction process, PaperRegister uses a large language model as the aggregating module $\mathcal{M}_{aggregate}$.
+Specifically, PaperRegister takes contents of sub-nodes as input and uses instructions to guide large language model in summarizing, condensing, and removing details to turn into summary text, thus obtaining content for upper-layer information node.
+The detailed process and instructions are in Appendix C.
+At this point, PaperRegister obtains node contents at various granular levels to compose the hierarchical register for $p^{(m)}$.
+And then merges register of all papers in corpus $\mathcal{C}$ to construct hierarchical index tree $\mathcal{I}_{h}$, as shown in following formula:
+$\mathcal{I}_{h}=\{\{{\rm I}_{i,j}\}_{j=1}^{Z}\}_{i=1}^{L}=\{\{\mathcal{M}_{idx}\{c_{i,j}^{(m)}\}_{m=1}^{|\mathcal{C}|}\}_{j=1}^{Z}\}_{i=1}^{L}$
+where $\mathcal{M}_{idx}$ is the indexing module such as BM25 and DPR, each ${\rm I}_{i,j}$ is an index in hierarchical index tree, corresponding to a kind of view for corpus.
+2.3 Online Adaptive Retrieval
+Based on the offline hierarchical indexing above, PaperRegister perform online adaptive retrieval via suitable indexes from hierarchical index tree.
+Specifically, PaperRegister first identifies views involved in the query, determining corresponding indexes from hierarchical index tree, then conduct retrieval by matching query against these indexes.
 View Identifying.
-In order to achieve adaptive retrieval, PaperRegister first uses a view recognizer to identify the view $v_{k}$ in the input query $q$, represented as follows:
+To achieve adaptive retrieval, PaperRegister first uses a view recognizer to identify the view $v_{k}$ in query $q$, represented as follows:
 $\{{v}_{k}\}_{k=1}^{K}=\mathcal{M}_{identify}(q)$ (6)
 where the candidate set of ${v}_{k}$ is all node paths in the hierarchical register schema, $\mathcal{M}_{identify}$ is employed base on a small-scale language model with special training, which will be explained in next section.
-And to ensure that identifying results can cover the real views of query as more as possible, $\mathcal{M}_{identify}$ uses the beam search strategy (Holtzman et al.
-2020) to sample the top-$K$ output views.
+And to ensure that identifying results can cover the real views of query as more as possible, $\mathcal{M}_{identify}$ uses the beam search strategy Holtzman2020The to sample the top-$K$ output views.
 View-based Matching.
-After identifying views in the input query, PaperRegister looks up corresponding indexes from the hierarchical index tree $\mathcal{I}_{h}$, as following:
+After identifying views in query, PaperRegister looks up corresponding indexes from hierarchical index tree $\mathcal{I}_{h}$, as follows:
 $\{{\rm I}_{k}\}_{k=1}^{K}=\mathcal{M}_{lookup}(\mathcal{I}_{h},\{{v}_{k}\}_{k=1}^{K})$ (7)
-where each ${\rm I}_{k}$ is consisted of $\{c_{k}^{(m)}\}_{m=1}^{|\mathcal{C}|}$ as in Formula 5.
-Then these indexes are used to calculate relevance score of input query $q$ and each paper $p^{(m)}$, as following:
+where ${\rm I}_{k}$ is consisted of $\{c_{k}^{(m)}\}_{m=1}^{|\mathcal{C}|}$ as Formula 5.
+Then these indexes are used to calculate relevance score of input query $q$ and each paper $p^{(m)}$  :
 $s(q,p^{(m)})=\max\{\mathcal{M}_{rel}(q,~{c}_{k}^{(m)})_{k=1}^{K}\}$ (8)
-where $\mathcal{M}_{rel}$ is relevance module such as BM25 and DPR.
-Finally, the retrieved papers with top-$M$ relevance score, $\{p^{(m)}\}_{m=1}^{M}$, are selected as retrieval results for query $q$.
-View Recognizer Training
-For the PaperRegister system mentioned above, the view recognizer is a key aspect, which must both alleviate the latency and ensure the accuracy of view identifying.
-To this end, as shown in Figure 3, we use a small-scale language model with 0.6B parameters as the base model, followed by initial supervised fine-tuning (SFT), and then further enhance its view identifying capability by hierarchical-reward group relative policy optimization (GRPO) training.
+where $\mathcal{M}_{rel}$ is relevance module such as BM25.
+Finally, papers with top-$M$ relevance score, $\{p^{(m)}\}_{m=1}^{M}$, are selected as final results for query.
+3 View Recognizer Training
+For PaperRegister system, view recognizer is a key module, which must both alleviate latency and ensure accuracy.
+To this end, as in Figure 3, we use a small-scale language model with 0.6B parameters as base, first do supervised fine-tuning (SFT), and then further enhance capability by hierarchical-reward group relative policy optimization (GRPO).
 Training Data.
@@
 $\mathcal{D}_{train}=\{q_{j},v_{j}\}_{j=1}^{|\mathcal{D}_{train}|}$ (9)
-where $q_{j}$ is a query, $v_{j}$ is the golden view of query, which is a node path in hierarchical register schema.
-To ensure the stability of training, we make the approximate same number of each view in the training dataset through manual sampling.
+where $q_{j}$ is a query, $v_{j}$ is golden view of query, which is node path in hierarchical register schema.
 Supervised Fine-tuning.
-To give the small-scale view recognizer basic identifying ability and make it easier for subsequent reinforcement learning training, we first perform SFT by minimizing the following loss:
-$\mathcal{L}_{SFT}(\theta)=-\mathbb{E}_{(q,v)\sim\mathcal{D}_{train}}\sum_{t=1}^{|v|}\log\pi_{\theta}(v_{t}|q,v_{<t})$ (10)
-where $\pi_{\theta}$ is the model and $v_{t}$ is the $t$-th token of golden view.
+To give small-scale view recognizer basic identifying ability and make it easier for subsequent reinforcement learning, we first perform SFT by minimizing following loss:
+$\mathcal{L}_{SFT}(\theta)=-\mathbb{E}_{(q,v)\sim\mathcal{D}_{train}}\sum_{t=1}^{|v|}\log\pi_{\theta}(v_{t}|q,v_{<t})$
+where $\pi_{\theta}$ is model, $v_{t}$ is the $t$-th golden view token.
 Hierarchical-reward GRPO.
-To further strengthen the view identifying ability of small-scale view recognizer, we conduct GRPO (Shao et al.
-2024) training with hierarchical reward, which is calculated based on the node path’s closeness level of predicted view and golden view in the hierarchical register schema.
-Specially, considering the hierarchical feature of register schema, different wrong predicted views should not be treated equally.
-For example, if the golden view’s path in the hierarchical register schema is Abstract-Method-Implementation-Operation, and two predicted views are Abstract-Method-Implementation-Module and Abstract-Experiment-Dataset, although both are wrong, the first one is better than the second one because it is closer to golden view in the hierarchical register schema, suggesting a higher information overlap with the golden view.
-Based on this idea, we measure the path overlap between the predicted view $\hat{v}_{j}$ and the golden view $v_{j}$ in the hierarchical register schema, obtaining the hierarchical reward:
+To strengthen the capability of small-scale view recognizer, we conduct GRPO Shao2024DeepSeekMathPT with hierarchical reward, which is calculated via the closeness level of predicted view and golden view in hierarchical register schema.
+Specially, considering the hierarchical-dependency feature of register schema, different wrong predicted views should not be treated equally.
+For example, the golden view of query is represented as Abstract-Method-Implementation-Operation in schema, and two predicted views are Abstract-Method-Implementation-Module and Abstract-Experiment-Dataset respectively.
+Although both are wrong, the first one is better than the second because it is closer to golden view in hierarchical register schema, suggesting a higher information overlap with golden view.
+Based on this, we measure path overlap between predicted view $\hat{v}_{j}$ and golden view $v_{j}$ in hierarchical register schema, obtaining hierarchical reward:
 $r=\frac{\mathcal{M}_{overlap}(v_{j},\hat{v}_{j})}{|\hat{v}_{j}|}+\frac{\mathcal{M}_{overlap}(v_{j},\hat{v}_{j})}{|v_{j}|}$ (11)
-where $\mathcal{M}_{overlap}$ returns the overlapping level of two inputs, $|\hat{v}_{j}|$ means the path length of predicted view in the hierarchical register schema and $|v_{j}|$ is that of golden view.
-Then this hierarchical reward is used in GRPO training, which is by maximizing the following objective:
-$\mathcal{J}_{GRPO}(\theta)={}\mathbb{E}_{\left[q\sim D_{train},\left\{v_{i}\right\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(v\mid q)\right]}\frac{1}{G}\sum_{i=1}^{G}\frac{1}{\left|v_{i}\right|}\sum_{t=1}^{\left|v_{i}\right|}\Biggl{\{}$
-$\min\Biggl{[}\frac{\pi_{\theta}\left(v_{i,t}\mid q,v_{i,<t}\right)}{\pi_{\theta_{\text{old}}}\left(v_{i,t}\mid q,v_{i,<t}\right)}\hat{A}_{i,t},\text{clip}(\frac{\pi_{\theta}\left(v_{i,t}\mid q,v_{i,<t}\right)}{\pi_{\theta_{\text{old}}}\left(v_{i,t}\mid q,v_{i,<t}\right)},1-\epsilon,1+\epsilon)\hat{A}_{i,t}\Biggr{]}$
-$-\beta\mathbb{D}_{\text{KL}}\left[\pi_{\theta}\parallel\pi_{\text{ref}}\right]\Biggr{\}},~\text{where}~\hat{A}_{i,t}=\frac{r_{i}-mean(\{r_{1},r_{2},...,r_{G}\})}{std(\{r_{1},r_{2},...,r_{G}\})}$
-where $\pi_{\theta_{\text{old}}}$ is initialized by the SFT model, $G$ is the number of each group, $\epsilon$ and $\beta$ are two hyper-parameters.
-Method BM25-based Matching DPR-based Matching
+where $\mathcal{M}_{overlap}$ returns overlap level of inputs, $|\hat{v}_{j}|$ means path length of predicted view in hierarchical register schema, and $|v_{j}|$ is that of golden view.
+Then this hierarchical reward is used in GRPO, which is by maximizing the following objective:
+$\begin{aligned} \mathcal{J}_{GRPO}(\theta)={}&\mathbb{E}_{\left[q\sim D_{train},\left\{v_{i}\right\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(v\mid q)\right]}\frac{1}{G}\sum_{i=1}^{G}\frac{1}{\left|v_{i}\right|}\sum_{t=1}^{\left|v_{i}\right|}\Biggl\{\\ &\min\Biggl[\frac{\pi_{\theta}\left(v_{i,t}\mid q,v_{i,<t}\right)}{\pi_{\theta_{\text{old}}}\left(v_{i,t}\mid q,v_{i,<t}\right)}\hat{A}_{i,t},\text{clip}(\frac{\pi_{\theta}\left(v_{i,t}\mid q,v_{i,<t}\right)}{\pi_{\theta_{\text{old}}}\left(v_{i,t}\mid q,v_{i,<t}\right)},1-\epsilon,1+\epsilon)\hat{A}_{i,t}\Biggr]\\ &-\beta\mathbb{D}_{\text{KL}}\left[\pi_{\theta}\parallel\pi_{\text{ref}}\right]\Biggr\},~\text{where}~\hat{A}_{i,t}=\frac{r_{i}-mean(\{r_{1},r_{2},...,r_{G}\})}{std(\{r_{1},r_{2},...,r_{G}\})}\end{aligned}$
+where $\pi_{\theta_{\text{old}}}$ is initialized by SFT model, $G$ is number of each group, $\epsilon$ and $\beta$ are hyper-parameters.
+Method BM25-based Paper Search DPR-based Paper Search
 LitSearch F.g.Search-1 F.g.Search-2 F.g.Search-3 LitSearch F.g.Search-1 F.g.Search-2 F.g.Search-3
@@
 Total Paper 64.2 68.9 79.7 82.4 81.4 83.5 84.2 86.2 74.8 81.9 72.9 78.7 68.8 73.8 65.8 71.0
-Query Paraphrasing
-Rewriting 61.9 69.6 58.4 64.2 54.1 59.8 50.9 56.3 76.0 83.2 67.1 72.6 60.4 65.2 55.2 60.9
-HyDE 68.9 75.4 60.3 66.3 54.3 60.0 51.5 57.6 76.5 83.4 66.8 72.3 58.9 65.7 56.3 62.4
-CSQE 69.0 73.8 59.1 64.4 54.2 58.6 51.3 55.9 77.9 81.7 65.8 71.1 60.2 65.5 55.2 60.9
-Paper Splitting
-$\text{Chunk}_{avg}$ 58.1 67.6 68.1 74.2 66.6 71.9 65.9 71.6 58.1 67.6 51.1 58.3 46.4 53.4 46.4 52.2
-$\text{Chunk}_{max}$ 67.9 75.5 80.0 83.1 81.4 84.7 85.3 87.6 67.9 75.5 79.8 83.1 76.6 80.5 79.0 82.1
-$\text{Paragraph}_{avg}$ 29.7 38.3 58.9 66.6 59.8 66.5 58.2 66.2 20.5 22.8 23.3 30.6 23.1 29.8 22.4 29.2
-$\text{Paragraph}_{max}$ 64.3 70.9 73.8 78.8 75.4 79.9 80.8 83.9 79.5 85.0 79.2 82.8 76.5 81.4 78.8 81.8
+Query Enhancing
+Rewriting ma2023query 61.9 69.6 58.4 64.2 54.1 59.8 50.9 56.3 76.0 83.2 67.1 72.6 60.4 65.2 55.2 60.9
+HyDE gao2023precise 68.9 75.4 60.3 66.3 54.3 60.0 51.5 57.6 76.5 83.4 66.8 72.3 58.9 65.7 56.3 62.4
+CSQE lei2024corpus 69.0 73.8 59.1 64.4 54.2 58.6 51.3 55.9 77.9 81.7 65.8 71.1 60.2 65.5 55.2 60.9
+Multi-field Indexing
+$\text{Chunk}_{avg}$ li2025multifield 58.1 67.6 68.1 74.2 66.6 71.9 65.9 71.6 49.7 58.6 51.1 58.3 46.4 53.4 46.4 52.2
+$\text{Chunk}_{max}$ li2025multifield 67.9 75.5 80.0 83.1 81.4 84.7 85.3 87.6 72.6 79.7 79.8 83.1 76.6 80.5 79.0 82.1
+$\text{Paragraph}_{avg}$ li2025multifield 29.7 38.3 58.9 66.6 59.8 66.5 58.2 66.2 20.5 22.8 23.3 30.6 23.1 29.8 22.4 29.2
+$\text{Paragraph}_{max}$ li2025multifield 64.3 70.9 73.8 78.8 75.4 79.9 80.8 83.9 79.5 85.0 79.2 82.8 76.5 81.4 78.8 81.8
 Hierarchical Register Indexing
-PaperRegister (Ours) 69.7 76.4 89.7 90.9 88.0 89.0 87.5 88.7 81.0 87.1 84.1 87.1 79.9 82.5 80.8 82.9
-Table 1: Main experimental results on paper-searching tasks across various granularity, where the granularity is coarse-to-fine from LitSearch to F.g.Search-3.
-The table shows that PaperRegister is a powerful solution to addressing flexible-grained paper search and the advantages of PaperRegister become more pronounced at finer granular queries.
-Experiments
-Experimental Settings
+PaperRegister (Ours) 69.7 76.4 89.7* 90.9* 88.0* 89.0* 87.5* 88.7* 81.0* 87.1* 84.1* 87.1* 79.9* 82.5* 80.8* 82.9*
+Table 1: Main results on paper search across various granularity, where granularity is coarse-to-fine from LitSearch to F.g.Search-3.
+Results shows that PaperRegister is an effective system for flexible-grained paper search and the advantage of PaperRegister become more pronounced at finer granular tasks.
+The * indicates statistical significance (p < 0.05) compared with the best baseline.
+And experiment results by more metrics are reported in Appendix E.
+4 Experiments
+4.1 Experimental Settings
 Datasets.
-We first use LitSearch (Ajith et al.
-2024), mainly containing coarse-grained queries.
-And then we build a new dataset, Flexible-grained Search, covering queries across various granularity and including data for test, development, and training.
-The detailed statistical information of dataset is shown in Appendix D.
-And the metrics used and reported in experiments are R@5 (recall@5) and R@10 (recall@10).
-(1) LitSearch (Ajith et al.
-2024).
-By using GPT-4 to rewrite citations in papers and asking authors to write, this dataset constructs 597 paper search queries, mainly involving coarse-grained topics like “Where can I find research on evaluating consistency in generated summaries”.
+To conduct experiments for flexible-grained paper search, we first use LitSearch ajith2024litsearch, which mainly contains coarse-grained queries, and then we build Flexible-grained Search, a new dataset covering queries across various granularity, including data for test, development and training.
+The metrics reported in experiments are R@5 (recall@5) and R@10 (recall@10).
+And detailed statistical information is in the Appendix D.
+(1) LitSearch ajith2024litsearch.
+By using GPT-4 to rewrite citations in papers and asking authors to write, this dataset constructs 597 paper search queries, mainly involving coarse-grained topics like “Where can I find some researches on evaluating consistency in generated summaries”.
 (2) Flexible-grained Search.
-We collect 4,200 papers in machine learning field as corpus and perform the offline hierarchical indexing.
-To build test data across various granularity, we pick hierarchical register of 2,100 papers, employ Qwen3-32B to find original paper texts related to each content in register, then use Qwen3-32B to generate queries based on these texts, obtaining 5,644 test data including 3 parts: general-granularity (F.g.Search-1), fine-granularity (F.g.Search-2), and very fine-granularity (F.g.Search-3).
-The format of each test data is like $\{q_{j},p_{j}\}$.
-To build training and development data, we use register of another 2,100 paper and directly generate queries for contents in each paper register, where each query is with a golden view, the format of training and development data is like $\{q_{j},v_{j},p_{j}\}$.
-Then we do random splitting and get 13,824 training data and 695 development data, the training data are used in view recognizer training part as illustrated in the Formula 9 and development data are used for analysis in the Table 3.
+We collect 4,200 papers in the arXiv platform as corpus and perform the offline hierarchical indexing.
+To build training and development data, we pick 2,100 paper registers, and generate query based on each node content in each register by LLM, where the node path is golden view of query.
+Then we randomly split, getting 13,824 training data and 695 development data, training data are used in view recognizer training, and development data are for analysis in Table 3.
+The format of training and development data is $\{q_{j},v_{j},p_{j}\}$.
+To build test data across various granularity, we pick the other 2,100 paper registers.
+And in order to prevent data leakage, we do not directly use node content to generate query.
+Instead, we employ LLM to find original paper text related to each node, then generate queries based on these text, obtaining 5,644 test data.
+Thanks to the hierarchical feature of register, test data cover: F.g.Search-1 (general granularity), F.g.Search-2 (fine granularity), and F.g.Search-3 (very fine granularity).
+Format of each test data is $\{q_{j},p_{j}\}$, and LLM in this section is Qwen3-32B qwen3.
 Selected Baselines.
-We select baselines from commonly used or recent methods suitable for paper search tasks.
-(1) Direct Matching. These methods directly use paper titles, abstracts or total text of papers to construct index of corpus, and retrieval by matching query with these contents.
-(2) Query-paraphrasing Methods.
-These methods aim to improve paper‐searching performance by paraphrasing the input query, including: Rewriting (Ma et al.
-2023), which uses a LLM to rewrite original query.
-HyDE (Gao et al.
-2023), which uses a LLM to generate a fake document based on input query and retrieves real documents by this fake document.
-CSQE (Lei et al.
-2024), which initially retrieves several documents and then uses a LLM to expand original query based on these initially-retrieved documents.
-We use Qwen3-32B as the LLM when conducting these baselines.
-(3) Paper-splitting Methods.
-These methods split the original paper into multiple short parts, then calculate the similarity between the query and each short part, and finally integrate all the similarity to determine the overall similarity of query and original paper.
-In experiments, we employ four kinds of settings, including splitting the original paper to the fixed length of 512-token chunk or by the raw paragraph, taking the average similarity or the maximum similarity of all parts as the final score, represented as $\text{Chunk}_{avg}$, $\text{Chunk}_{max}$, $\text{Paragraph}_{avg}$, and $\text{Paragraph}_{max}$, respectively.
-Method BM25-based Matching DPR-based Matching
+We select baselines from commonly used or advanced methods suitable for paper search tasks.
+(1) Direct Matching.
+These methods directly use the paper title or abstract to construct corpus index, and then perform retrieval by matching query with these contents.
+In addition, to demonstrate that the effectiveness of PaperRegister does not simply stem from using total paper text, we include a method that constructs the corpus index using the total paper text into this category of baselines.
+(2) Query Enhancing.
+These methods aim to improve paper search performance by paraphrasing the input query, including: Rewriting ma2023query, which uses a LLM to rewrite original query.
+HyDE gao2023precise, which uses a LLM to generate a fake document based on input query and retrieves real documents by this fake document.
+CSQE lei2024corpus, which initially retrieves several documents and then uses a LLM to expand original query based on these initially-retrieved documents.
+We use Qwen3-32B as LLM for these baselines.
+(3) Multi-field Indexing.
+These methods split original document into multiple parts, get flat multi-field index, then calculate similarity by selected indexes and integrate to determine overall relevance li2025multifield.
+In experiments, due to the lack of a universal multiple-filed partitioning for all kinds of paper, we employ four settings, including splitting the original paper to fixed length of 512-token chunk or by raw paragraph, taking average similarity or maximum similarity of all parts as final score, represented as $\text{Chunk}_{avg}$, $\text{Chunk}_{max}$, $\text{Paragraph}_{avg}$, and $\text{Paragraph}_{max}$, respectively.
+Implementation Details.
+In the offline stage, we employ Qwen3-32B qwen3 as the large language model in process of fine-grained content extracting and bottom-up content aggregating, and deploy it as API by vllm [^1] https://pypi.org/project/vllm/ on two A-100 80G GPU for convenience.
+In the online stage, we set $K$ as 5 and $M$ as 5 or 10, use Qwen3-0.6B qwen3 as the base model of view recognizer, employ a prefix tree-based restricted decoding strategy tang2024selfretrieval in view identifying process to prevent the model outputting irrelevant token, and conduct view-based matching process via rank-bm25222https://pypi.org/project/rank-bm25/ for BM25 and gte-Qwen2-7B-instruct li2023towards for DPR.
+For the view recognizer training, we use TRL [^3] https://github.com/huggingface/trl framework to conduct SFT and GRPO, with train epoch as 5 in SFT, $G$ as 5 and train epoch as 2 in GRPO, and all other parameters as default value.
+Method BM25-based Paper Search DPR-based Paper Search
 LitSearch F.g.Search-1 F.g.Search-2 F.g.Search-3 LitSearch F.g.Search-1 F.g.Search-2 F.g.Search-3
 R@5 R@10 R@5 R@10 R@5 R@10 R@5 R@10 R@5 R@10 R@5 R@10 R@5 R@10 R@5 R@10
-PaperRegister (Ours) 69.7 76.4 89.7 90.9 88.0 89.0 87.5 88.7 81.0 87.1 84.1 87.1 79.9 82.5 80.8 82.9
+PaperRegister 69.7 76.4 89.7 90.9 88.0 89.0 87.5 88.7 81.0 87.1 84.1 87.1 79.9 82.5 80.8 82.9
 w/ only layer-1 64.5 70.4 87.8 90.6 77.4 80.2 73.5 76.7 79.0 83.5 82.5 85.2 73.4 77.1 68.1 72.2
@@
 w/ only layer-3 64.6 71.7 79.9 81.8 83.7 85.0 86.8 88.1 78.8 85.2 77.8 81.1 77.3 80.0 80.4 82.6
-Table 2: Ablation results for the hierarchical register indexing.
-The table shows that hierarchical index tree are important for flexible-grained paper search and different layers in the hierarchical register schema serve queries at different granularity.
-Implementation Details.
-In the offline stage, we employ Qwen3-32B (Qwen 2025) as the large language model in process of fine-grained content extracting and bottom-up content aggregating, and deploy it as API by vllm [^1] https://pypi.org/project/vllm/ on two A-100 80G GPU for convenience.
-In the online stage, we set $K$ as 5 and $M$ as 5 or 10, use Qwen3-0.6B (Qwen 2025) as the base model of view recognizer, employ a prefix tree-based restricted decoding strategy (Tang et al.
-2024) in the view identifying process to prevent the model outputting irrelevant token, and conduct the view-based matching process via rank-bm25222https://pypi.org/project/rank-bm25/ for BM25 and gte-Qwen2-7B-instruct (Li et al.
-2023) for DPR.
-For the view recognizer training, we use TRL [^3] https://github.com/huggingface/trl framework to conduct SFT and GRPO.
-We set train epoch as 5 in SFT, $G$ as 5 and train epoch as 2 in GRPO, and keep all other parameters as default value in TRL.
-When conducting baselines, we keep the same settings as above.
-Overall Results
-Results compared with baselines are shown in the Table 1, there are two main conclusions based on the results:
-(1) PaperRegister is a powerful solution to addressing flexible-grained paper search tasks.
+Table 2: Ablation for hierarchical register indexing.
+Results show that hierarchical index tree is important for flexible-grained paper search and different layers in hierarchical register schema serve queries at different granularity.
+4.2 Overall Results
+Results compared with baselines are shown in the Table 1, there are two main conclusions:
+(1) PaperRegister is an effective system to addressing flexible-grained paper search.
 As shown in the Table 1, PaperRegister demonstrates excellent performance in paper search tasks at various granularity, outperforming all baselines in both BM25-based matching and DPR-based matching settings, on both recall@5 and recall@10 metrics.
@@
 All in all, PaperRegister can improve a lot compared with various previous paper search methods.
-Figure 4: Paper search performance of the PaperRegister with different view recognizer. The figure shows strong recognizer is with obvious positive impact on overall system.
-(2) Compared to traditional methods, the advantages of PaperRegister become more pronounced at these finer granular queries.
-According to experimental setting, from LitSearch to F.g.Search-1 to F.g.Search-2 to F.g.Search-3, the granularity of the paper search query becomes increasingly finer.
-And the experimental results show that the performance improvement of PaperRegister over the method using abstract-based index becomes larger.
+(2) Compared to traditional methods, advantage of PaperRegister become more pronounced at finer granular queries.
+According to experimental setting, from LitSearch to F.g.Search-1 to F.g.Search-2 to F.g.Search-3, the granularity of query becomes increasingly finer.
+Experimental results show performance improvement of PaperRegister over abstract-based method becomes larger.
 For example, under the DPR-based matching setting, PaperRegister achieves recall@5 scores of 81.0, 84.1, 79.9, and 80.8 on the four datasets, respectively, while using the abstract-based index yields scores of 77.7, 69.4, 62.1, and 58.2.
 Here, PaperRegister delivers performance improvements of 3.3, 14.7, 17.8, and 22.6, respectively.
-Therefore, PaperRegister exhibits more improvements over baseline on finer-grained queries, which can to some extent validate rationality of this work’s motivation, that traditional paper search via abstract-based index struggles to handle fine-grained queries.
-Ablation Results
-To validate the importance of hierarchical register indexing, we employ ablation experiments by simplifying the hierarchical register schema.
+This can to some extent validate rationality of our motivation, traditional paper search methods via abstract-based index struggles to handle fine-grained queries.
+4.3 Ablation Results
+To validate importance of hierarchical register indexing, we employ ablation by simplifying register schema.
 As shown in Table 2, where “w/ only layer-1” means retaining only the coarsest-grained nodes, “w/ only layer-2” means retaining only the medium-grained nodes, and “w/ only layer-3” means retaining only the finest-grained nodes.
-Based on these simplified register schema for experiments, we get the following conclusions:
-(1) Hierarchical register indexing is important for ensuring the effectiveness of paper search.
-As shown in Table 2, compared to using the complete hierarchical register schema, using a register schema composed of nodes from any single layer leads to obvious performance drop.
+Base on the ablation results, we get the following conclusions:
+(1) Hierarchical register indexing is important for ensuring accurate paper search.
+Compared to using complete hierarchical register schema, using a schema composed of nodes from any single layer leads to obvious performance drop.
 For example, recall@5 of LitSearch under BM25-based matching drops from 69.7 to 64.5, 66.8, 64.6 for “w/ only layer-1”, “w/ only layer-2”, “w/ only layer-3”, respectively.
 And other datasets also show similar situation.
-Therefore, results strongly prove importance of hierarchical register indexing.
-(2) Different layers in the hierarchical register schema serve queries at different granularity.
+Therefore, results strongly prove the importance of hierarchical register indexing.
+Figure 4: Performance of PaperRegister with different view recognizer. The figure shows a strong recognizer is with obvious positive impact on the overall system.
+(2) Different layers in hierarchical register schema serve queries at different granularity.
 For example, under “w/ only layer-1” setting, the DPR-based matching and recall@5 metric of F.g.Search-1 drops from 84.1 to 82.5 (1.6 decrease), while F.g.Search-1 is from 80.8 to 68.1 (12.7 decrease), which is significantly larger than 1.6.
 Conversely, under “w/ only layer-3” setting, F.g.Search-1 drops from 84.1 to 77.8 (6.3 decrease), while F.g.Search-3 is from 80.8 to 80.4 (0.4 decrease), which is much smaller than 6.3.
-Therefore, nodes from different layers play specific roles in handling queries of different granularity, further strongly demonstrating necessity of hierarchical register indexing.
+Therefore, nodes of different layers play specific roles in handling queries of different granularity, further strongly proving necessity of hierarchical register indexing.
+4.4 Detailed Analysis
+In this section, we conduct detailed analysis to explore the compatibility and real-world practicality of PaperRegister. Due to the space constraint, we only report DPR-based performance for analysis.
+Effect of View Recognizer and Training.
+In order to strictly analyze the role of view recognizer in PaperRegister and the necessity of our training, we first examine relationship between the capability of view recognizer and the final performance of PaperRegister in Figure 4, then compare with Qwen3-32B and conduct ablation in Table 3.
+The detailed process and findings are in following:
+As shown in Figure 4, where “Bad Recognizer”,“Random Recognizer”,“Weak Recognizer” represent the completely incorrect, randomly predicting, and weak-performing view recognizers, respectively, the curve demonstrates a clear positive correlation between the capability of view recognizer and final performance of PaperRegister.
+Therefore, building an accurate view recognizer is a key factor for the PaperRegister system to achieve high performance in complex paper search.
 Recognizer ACC $\uparrow$ Latency (s) $\downarrow$
@@
 The table shows that the training process in this work is an effective approach to obtain a view recognizer with both high performance and low latency.
-Detailed Analysis
-In this section, we conduct detailed analysis including: the effect of view recognizer, the compatibility of PaperRegister with complex PaSa framework (He et al.
-2025), and an analysis of the online search efficiency.
-Due to the space constraint, we only report the performance that is based on DPR-based matching for analysis in this section.
-Effect of View Recognizer and Training.
-To analyze the role of view recognizer in PaperRegister and necessity of our training, we first examine relationship between recognizer capability and PaperRegister performance in Figure 4, then compare with Qwen3-32B and conduct ablation in Table 3.
-As shown in Figure 4, where “Bad Recognizer”,“Random Recognizer”,“Weak Recognizer” represent the completely incorrect, randomly predicting, and weak-performing view recognizers, respectively, the curve demonstrates a clear positive correlation between the capability of view recognizer and final performance of PaperRegister.
-Therefore, building an accurate view recognizer is a key factor for the PaperRegister to achieve high-performance paper search.
 As shown in Table 3, We compare accuracy and latency of view identifying between the view recognizer in PaperRegister and Qwen3-32B (enable thinking) under zero-shot and few-shot settings.
@@
 Furthermore, we only keep SFT for the view recognizer training, or replace the hierarchical reward with direct 0-1 reward in GRPO training.
-Results show that the accuracy decreases to some extent.
-Therefore, training a small-scale language model with SFT and then hierarchical-reward GRPO is an effective approach to obtain the view recognizer with both high performance and low latency.
-Compatibility with Complex PaSa Framework.
-Considering that PaperRegister is mainly a retrieval module, and given that some recent complex paper search frameworks such as PaSa (He et al.
-2025) incorporate various modules like rewriting, retrieval, iteration, and filtering, in order to explore whether PaperRegister can be compatible with such complex frameworks and further enhance performance, we replace the original retrieval module in PaSa with PaperRegister.
-As shown in Figure 5, PaperRegister can further improve the performance of PaSa across paper search tasks at various granularity.
-Therefore, PaperRegister can be effectively adapted as a retrieval module into complex paper search frameworks and further enhance the capability.
-Figure 5: Performance of adding PaperRegister into the PaSa framework. The figure shows that PaperRegister can greatly cooperate with extra complex modules in PaSa framework.
-Analysis for Online Search Efficiency.
-Considering that the efficiency of online paper search system is a crucial aspect of its usability in practical scenarios, we compare the online search efficiency of PaperRegister with multiple baselines.
+Results show that accuracy decreases to some extent.
+Therefore, hierarchical-reward GRPO is effective to enhance the view recognizer achieving both high performance and low latency.
+Compatibility with PaSa.
+Given that some complex information retrieval frameworks such as PaSa he2025pasa incorporate various modules like rewriting, retrieval, iteration, and filtering, to explore whether PaperRegister can be compatible with such complex frameworks and further enhance paper search performance, we conduct experiment by replacing the original search module in PaSa with PaperRegister.
+As shown in Figure 5, PaperRegister can further improve performance of PaSa on paper search tasks across various granularity.
+Therefore, PaperRegister can be effectively adapted as a search system into complex frameworks and further enhance the paper search capability.
+Figure 5: Performance of combining PaperRegister into PaSa framework. The figure shows that PaperRegister can greatly cooperate with complex modules in PaSa.
+Online Search Efficiency.
+Considering that the efficiency of online paper search system is a very crucial aspect for ensuring its usability in real-world scenarios, we compare the online search efficiency of PaperRegister with multiple baselines.
 As shown in the Table 4, the online latency of these baselines is significantly higher than that of PaperRegister, which limits the practical applicability of these methods in real-world scenarios.
-In contrast, PaperRegister demonstrates the acceptable online latency, proving its good potential for application in the real-world paper search.
-Related Work
-To achieve accurate paper search, previous related works mainly focus on enhancing the input query, with less attention given to the index of paper corpus.
-Some works improve paper search performance by rewriting the original query, such as using powerful large language models for rewriting based on some manually designed instructions (Ma et al.
-2023; Anand et al.
-2023), generating pseudo-documents to replace original query for retrieval real documents (Gao et al.
-2023; Li et al.
-2024a), or employing several powerful models as agent system to conduct multi-round rewriting and expansion (He et al.
-2025; Ren et al.
-2025).
-Another kind of works expands the input query through pseudo-relevance feedback, such as augmenting the query based on initially retrieved documents (Lei et al.
-2024; Li et al.
-2025a), or extracting several keywords from the corpus to enrich the information of original query (Kang et al.
-2024; Zhang et al.
-2025).
-Although these works can improve performance in regular paper search, they fail to address flexible-grained queries because they do not solve the flaws in the index of paper corpus.
-Recently, several works split original paper by fine-grained fields and then construct multiple-field indexes for retrieval (Sotaro et al.
-2024; Li et al.
-2025b; Shi et al.
-2025; Chen et al.
-2025).
-However, these works still cannot effectively handle flexible-grained queries because they only focus on a single granularity level.
-Compared with above related works, the largest advantage of PaperRegister is hierarchical register indexing and conducting paper search with hierarchical index tree, which can effectively handle flexible-grained paper search.
-Method Online Paper Search Latency (s) $\downarrow$
-Rewriting 9.3
-HyDE 20.7
-CSQE 33.5
-$\text{Chunk}_{max}$ 5.4
-$\text{Paragraph}_{max}$ 4.2
-PaperRegister (Ours) 2.5
-Table 4: Comparison of online latency among PaperRegister and several baselines. The table show that PaperRegister is with better real-world applicability with less latency.
-Conclusion
-In this work, noticed limitation of existing methods in paper search tasks, we propose PaperRegister, boosting flexible-grained paper search via hierarchical register indexing.
-In the offline stage, PaperRegister constructs a hierarchical index tree by fine-grained content extracting and bottom-up content aggregating.
-In the online stage, PaperRegister first identify views of input query, then matches the query against corresponding indexes from the hierarchical index tree.
-Furthermore, we construct the view recognizer by applying supervised finetuning and hierarchical-reward group relative policy optimization to a small-scale language model.
-Experiments on extensive flexible-grained paper search tasks demonstrate that PaperRegister is an effective solution, which reaches the SOTA performance, with improvement being more pronounced as query granularity finer.
-Therefore, this work offers a promising direction for developing more powerful paper search systems in future.
-Appendix A
-This part is about detailed illustration for the hierarchical register schema.
-Figure 6: The first kind of hierarchical register schema.
-Figure 7: The second kind of hierarchical register schema.
-Figure 8: The third kind of hierarchical register schema.
-Figure 9: The fourth kind of hierarchical register schema.
-Figure 10: The fifth kind of hierarchical register schema.
+In contrast, PaperRegister demonstrates relatively acceptable online latency, proving it has a good potential for application in the real-world paper search tasks.
+Reducing Indexing Time.
+Although PaperRegister achieves excellent performance, it requires a relatively long time for offline indexing.
+Therefore, we explore whether it can sacrifice a certain amount of performance to reduce time consumption.
+As shown in Table 5, we replace original two-step registering (extracting and aggregating) with directly extracting all contents.
+Results show this can significantly reduce the offline indexing time with acceptable performance degradation, which provides an alternative approach allowing a trade-off based on specific requirements in real-world application.
+5 Related Work
+Query Enhancing.
+Previous works mainly improve paper search by query enhancing, such as using LLM rewriting ma2023query; anand2023context, generating pseudo-documents to replace original query gao2023precise; li2024automir, employing powerful models as agent system for multi-round expansion he2025pasa; ren2025towards, augmenting via initially retrieved documents lei2024corpus; li2025pseudo, or extracting keywords from corpus to enrich query kang2024taxonomy; zhang2025scientific.
+Although these methods can improve regular paper search, they fail to address flexible-grained paper search because do not solve fundamental flaw of abstract-based index.
+Multi-filed Indexing.
+Recently, several studies improve by splitting documents into multiple parts to build a flat, multi-field index takeshita-etal-2024-gengo; li2025multifield; shi2025hypercube; chen2025enrichindex.
+However, these works are mainly limited to a single granularity level, still struggling to adapt to paper search that demands varying degrees of granularity.
+In contrast, PaperRegister fundamentally differs by introducing a hierarchical register index.
+Tree-structured organization contains information across various granularity, enabling system to conduct matching at the appropriate levels, thereby effectively supporting flexible-grained paper search.
+Method Avg. R@5 Avg. R@10 Online Latency (s) $\downarrow$
+Rewriting 64.7 70.5 9.3
+HyDE 64.6 70.9 20.7
+CSQE 64.8 69.8 33.5
+$\text{Chunk}_{max}$ 77.0 81.4 5.4
+PaperRegister (Ours) 81.5 84.9 2.5
+Table 4: Comparison of online latency. PaperRegister is with less latency for better real-world applicability.
+Method Avg. R@5 Avg. R@10 Indexing Time (h) $\downarrow$
+Abstract 66.9 72.1 /
+PaperRegister (Ours) 81.5 84.9 47.6
+w/o two-step registering 76.1 78.9 14.5
+Table 5: PaperRegister can support sacrificing certain performance to alleviate offline time consumption.
+6 Conclusion
+In this work, we propose PaperRegister, which can support paper search queries at flexible granularity.
+Specially, PaperRegister offline constructs a hierarchical index tree and online adaptively retrieve based on this tree.
+Furthermore, we design a powerful and low-latency view recognizer by applying supervised finetuning and hierarchical-reward group relative policy optimization.
+Experiments on flexible-grained paper search tasks demonstrate PaperRegister is an effective solution, with improvement being more pronounced as granularity finer.
+This work offers a promising direction for developing more powerful paper search systems in future.
+7 Limitations
+PaperRegister transforms raw paper into hierarchical paper register, thereby achieving strong performance in flexible-grained paper search.
+The main limitations are on the consumption associated with LLM utilization and risks stemming from the inherent limitations of LLM.
+Firstly, using LLM for offline hierarchical indexing requires substantial computational and storage resources.
+When the paper corpus is large, this may place high demands on computer hardware.
+Secondly, our experiments are conducted only on normal paper corpus, where LLM could effectively accomplish the specified task.
+We have not yet do validate under extreme conditions, such as with incomplete papers or papers from niche domains, primarily due to a lack of relevant experimental data.
+Therefore, mitigating computational resource consumption of PaperRegister, as well as verifying and improving effectiveness and robustness across a broader range of paper, remain important directions for the future work.
+Appendix A Detailed illustration for hierarchical register schema
 Since different types of papers have the different styles and formats, we design five kinds of hierarchical register schema, corresponding to five types of paper including algorithm innovation, benchmark construction, mechanism exploration, survey, and theory proof.
@@
 ==Abstract== abstract
-Appendix B
-This part is about detailed process and instructions for the fine-grained content extracting.
-Learning from several widely recognized works, in which large language models are proven to possess reliable capabilities for content extraction (Edge et al.
-2024; Li et al.
-2025c, 2024b, 2024c), PaperRegister uses a large language model as the extracting module.
+Appendix B Detailed process and instructions for fine-grained content extracting
+Learning from several widely recognized works, in which large language models are proven to possess reliable capabilities for content extraction edge2024local; li2025deepsolution; li2024structrag; li2024meta, PaperRegister uses a large language model as the extracting module.
 PaperRegister takes the node name and the text-formatted paper as input, uses instructions to guide in outputting corresponding content for the information node, and leaves it blank if the paper does not include corresponding content.
@@
 ==Schema== schema ==Paper== paper
-Appendix C
-This part is about detailed process and instructions for the bottom-up content extracting.
+Dataset Name # Corpus # Query
+LitSearch 3400 597
+F.g.Search-1 4200 1,922
+F.g.Search-2 4200 1,922
+F.g.Search-3 4200 1,800
+Table 6: Statistics of test data in experiments.
+Appendix C Detailed process and instructions for bottom-up content extracting
 PaperRegister takes contents of sub-nodes as input and uses instructions to guide the large language model in summarizing, condensing, and removing details to turn into summary text, thus obtaining the content for upper-layer information node.
@@
 ==Input Tree== tree
-Appendix D
-This part is about detailed statistical information of dataset for the experiments.
+Appendix D Detailed statistical information of dataset for experiments
 We first use LitSearch, mainly containing coarse-grained queries.
@@
 The numbers of queries included in F.g.Search-1, F.g.Search-2, and F.g.Search-3 are 1,922, 1,922, and 1,800, respectively.
+The statics is in Table 6.
+Appendix E Experimental results by more metrics
+To provide a more comprehensive illustration of the advancements of PaperRegister over baseline methods, we present experimental results in the Table 789 across six metrics: Precision@5, Precision@10, MAP@5, MAP@10, NDCG@5, and NDCG@10.
+The data consistently support the conclusions drawn in our main experimental tables, further confirming that PaperRegister is an effective system for addressing flexible-grained paper search and the advantages of PaperRegister become more pronounced when handling finer-grained queries.
+Figure 6: The first kind of hierarchical register schema (algorithm innovation).
+Figure 7: The second kind of hierarchical register schema (benchmark construction).
+Figure 8: The third kind of hierarchical register schema (mechanism exploration).
+Figure 9: The fourth kind of hierarchical register schema (survey and review).
+Figure 10: The fifth kind of hierarchical register schema (theory proof).
+Method BM25-based Paper Search DPR-based Paper Search
+LitSearch F.g.Search-1 F.g.Search-2 F.g.Search-3 LitSearch F.g.Search-1 F.g.Search-2 F.g.Search-3
+P@5 P@10 P@5 P@10 P@5 P@10 P@5 P@10 P@5 P@10 P@5 P@10 P@5 P@10 P@5 P@10
+Direct Matching
+Title 11.4 6.3 8.4 4.7 7.2 4.2 6.1 3.5 14.0 7.8 10.4 5.9 9.2 5.2 8.1 4.7
+Abstract 14.1 7.6 12.7 6.9 11.5 6.3 10.8 6.0 16.4 8.8 13.9 7.4 12.4 6.8 11.6 6.3
+Total Paper 13.5 7.3 15.9 8.2 16.3 8.4 16.8 8.6 16.0 8.8 14.6 7.9 13.8 7.4 13.2 7.1
+Query Enhancing
+Rewriting ma2023query 13.0 7.4 11.7 6.4 10.8 6.0 10.2 5.6 16.0 8.8 13.4 7.3 12.1 6.5 11.0 6.1
+HyDE gao2023precise 14.6 8.0 12.1 6.6 10.9 6.0 10.3 5.8 16.2 8.9 13.4 7.2 11.8 6.6 11.3 6.2
+CSQE lei2024corpus 14.4 7.8 11.8 6.4 10.8 5.9 10.3 5.6 16.5 8.7 13.2 7.1 12.0 6.5 11.0 6.1
+Multi-field Indexing
+$\text{Chunk}_{avg}$ li2025multifield 12.2 7.2 13.6 7.4 13.3 7.2 13.2 7.2 10.5 6.2 10.2 5.8 9.3 5.3 9.3 5.2
+$\text{Chunk}_{max}$ li2025multifield 14.3 7.9 16.0 8.3 16.3 8.5 17.1 8.8 15.2 8.5 16.0 8.3 15.3 8.1 15.8 8.2
+$\text{Paragraph}_{avg}$ li2025multifield 6.2 4.1 11.8 6.7 12.0 6.7 11.6 6.6 4.3 2.4 4.7 3.1 4.6 3.0 4.5 2.9
+$\text{Paragraph}_{max}$ li2025multifield 13.5 7.5 14.8 7.9 15.1 8.0 16.2 8.4 16.7 9.0 15.8 8.3 15.3 8.1 15.8 8.2
+Hierarchical Register Indexing
+PaperRegister (Ours) 14.6 8.0 17.9 9.1 17.6 8.9 17.5 8.9 17.1 9.3 16.8 8.7 16.0 8.3 16.2 8.3
+Table 7: Precision@5 and Precision@10 results on paper search across various granularity.
+Method BM25-based Paper Search DPR-based Paper Search
+LitSearch F.g.Search-1 F.g.Search-2 F.g.Search-3 LitSearch F.g.Search-1 F.g.Search-2 F.g.Search-3
+M@5 M@10 M@5 M@10 M@5 M@10 M@5 M@10 M@5 M@10 M@5 M@10 M@5 M@10 M@5 M@10
+Direct Matching
+Title 42.2 42.9 33.4 34.1 28.8 29.6 23.5 24.2 54.9 55.9 41.3 42.2 36.3 37.2 32.2 33.0
+Abstract 52.4 53.0 54.9 55.6 49.4 50.1 45.8 46.5 65.1 65.9 59.8 60.5 51.9 52.7 47.2 47.8
+Total Paper 51.7 52.4 74.3 74.6 76.1 76.4 79.1 79.3 60.5 61.6 63.2 64.0 59.2 59.9 55.6 56.3
+Query Enhancing
+Rewriting ma2023query 47.2 48.3 48.7 49.5 44.8 45.5 41.2 41.9 62.3 63.4 56.7 57.4 50.1 50.8 44.8 45.6
+HyDE gao2023precise 54.1 55.1 50.1 50.9 44.6 45.3 41.7 42.6 64.3 65.3 56.4 57.1 48.7 49.6 45.1 45.9
+CSQE lei2024corpus 54.1 54.8 41.2 42.0 37.8 38.4 36.7 37.4 65.2 65.8 54.1 54.9 48.4 49.1 44.4 45.1
+Multi-field Indexing
+$\text{Chunk}_{avg}$ li2025multifield 40.6 42.0 54.4 55.2 54.9 55.6 54.4 55.2 34.9 36.1 39.4 40.4 36.6 37.5 35.1 35.9
+$\text{Chunk}_{max}$ li2025multifield 56.1 57.1 73.6 74.0 75.3 75.8 79.5 79.8 62.0 63.0 71.7 72.2 68.4 68.9 69.4 69.9
+$\text{Paragraph}_{avg}$ li2025multifield 19.0 20.2 44.9 46.0 46.2 47.1 44.8 45.9 15.0 15.3 15.9 16.8 15.7 16.6 15.1 15.9
+$\text{Paragraph}_{max}$ li2025multifield 50.7 51.7 66.4 67.0 67.2 67.8 73.2 73.6 66.1 67.0 70.8 71.3 68.5 69.1 68.7 69.1
+Hierarchical Register Indexing
+PaperRegister (Ours) 57.4 58.3 86.2 86.4 84.9 85.0 85.3 85.5 70.0 70.9 77.4 77.8 73.7 74.0 72.9 73.1
+Table 8: MAP@5 and MAP@10 results on paper search across various granularity.
+Method BM25-based Paper Search DPR-based Paper Search
+LitSearch F.g.Search-1 F.g.Search-2 F.g.Search-3 LitSearch F.g.Search-1 F.g.Search-2 F.g.Search-3
+N@5 N@10 N@5 N@10 N@5 N@10 N@5 N@10 N@5 N@10 N@5 N@10 N@5 N@10 N@5 N@10
+Direct Matching
+Title 45.4 47.1 35.6 37.3 30.6 32.5 25.2 26.8 58.2 60.3 44.0 46.2 38.7 40.7 34.3 36.4
+Abstract 56.3 57.7 57.1 58.8 51.4 53.1 47.9 49.6 68.5 70.3 62.2 63.8 54.5 56.4 50.0 51.5
+Total Paper 55.0 56.6 75.6 76.5 77.5 78.1 80.3 81.0 64.4 66.8 65.7 67.5 61.6 63.3 58.2 59.8
+Query Enhancing
+Rewriting ma2023query 51.2 53.7 51.1 53.0 47.1 48.9 43.6 45.4 66.0 68.4 59.3 61.1 52.7 54.3 47.4 49.3
+HyDE gao2023precise 58.1 60.3 52.7 54.6 47.0 48.9 44.2 46.1 67.6 69.9 59.0 60.8 51.3 53.5 47.9 49.8
+CSQE lei2024corpus 58.0 59.7 45.7 47.4 41.9 43.4 40.4 41.9 68.6 69.8 57.1 58.8 51.4 53.0 47.1 48.9
+Multi-field Indexing
+$\text{Chunk}_{avg}$ li2025multifield 45.2 48.4 57.8 59.8 57.8 59.5 57.3 59.1 38.8 41.8 42.3 44.7 39.1 41.3 37.9 39.8
+$\text{Chunk}_{max}$ li2025multifield 59.2 61.7 75.2 76.2 76.8 77.9 81.0 81.7 64.9 67.3 73.8 74.8 70.5 71.7 71.8 72.8
+$\text{Paragraph}_{avg}$ li2025multifield 21.8 24.7 48.4 50.9 49.6 51.7 48.2 50.8 16.4 17.2 17.7 20.0 17.5 19.7 16.9 19.0
+$\text{Paragraph}_{max}$ li2025multifield 54.4 56.6 68.2 69.9 69.3 70.7 75.1 76.1 69.7 71.5 72.9 74.1 70.5 72.1 71.2 72.2
+Hierarchical Register Indexing
+PaperRegister (Ours) 60.7 62.9 87.1 87.5 85.7 86.0 85.9 86.3 73.0 75.0 79.1 80.1 75.2 76.1 74.9 75.5
+Table 9: NDCG@5 and NDCG@10 results on paper search across various granularity.